### v1

# A semiconcavity approach to stability of entropic plans and exponential convergence of Sinkhorn’s algorithm

Abstract.
We study stability of optimizers and convergence of Sinkhorn’s algorithm in the framework of entropic optimal transport. We show entropic stability for optimal plans in terms of the Wasserstein distance between their marginals under a semiconcavity assumption on the sum of the cost and one of the two entropic potentials. When employed in the analysis of Sinkhorn’s algorithm, this result gives a natural sufficient condition for its exponential convergence, which does not require the ground cost to be bounded. By controlling from above the Hessians of Sinkhorn potentials in examples of interest, we obtain new exponential convergence results. For instance, for the first time we obtain exponential convergence for log-concave marginals and quadratic costs for all values of the regularization parameter. Moreover, the convergence rate has a linear dependence on the regularization: this behavior is sharp and had only been previously obtained for compact distributions [13]. Other interesting new applications include subspace elastic costs [16], weakly log-concave marginals, marginals with light tails, where, under reinforced assumptions, we manage to improve the rates obtained in [28], the case of unbounded Lipschitz costs, and compact Riemannian manifolds.
While this work was written, AC was associated to INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”) and the group GNAMPA
GG is supported by the PRIN project GRAFIA (CUP: E53D23005530006) and MUR Departement of Excellence Programme 2023-2027 MatMod@Tov (CUP: E83C23000330006). GG is grateful to the Department of Mathematics “Tullio Levi-Civita” for its hospitality, where part of this work has been carried out.
While this work was written, LT was associated to INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”) and the group GNAMPA
Mathematics Subject Classification (2020)
49Q22, 68Q87, 68W40, 60E15, 90C25
1. Introduction
Given two Polish spaces $\mathcal{X},\mathcal{Y}$, marginal probability distributions $\rho\in\mathcal{P}(\mathcal{X}),\nu\in\mathcal{P}(\mathcal{Y})$, and a cost function $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, the entropic optimal transport problem (EOT) reads as
(EOT) $\inf_{\pi\in\Pi(\rho,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c\,\mathrm{d}\pi+\varepsilon\,\mathrm{KL}(\pi|\rho\otimes\nu),$
where $\Pi(\mu,\nu)$ is the set of couplings of $\mu$ and $\nu$, $\mathrm{KL}$ denotes the Kullback-Leibler divergence (also known as relative entropy), and $\varepsilon>0$ is a regularization parameter. The study of EOT has greatly intensified since the observation [22] that adding an entropic penalty in the objective function of the Monge-Kantorovich problem (corresponding to $\varepsilon=0$ in (EOT)) leads to a more convex, more regular, and numerically more tractable optimization task, thus opening new perspectives for the computation of transport distances in machine learning and beyond, see [49]. Much of the success of entropic regularization techniques in applications can be attributed to the fact that EOT can be solved by means of an exponentially-fast matrix scaling algorithm, Sinkhorn’s algorithm, and to the fact that EOT is more stable than the Monge-Kantorovich problem with respect to variations in the cost or marginals. Because of this, considerable efforts have been made over the last decade to turn these intuitions into sound mathematical statements. This has produced many important contributions which we shall discuss in more detail below, nonetheless, several open questions remain. For example, exponential convergence of Sinkhorn algorithm is not well understood when both the marginals’ support and the ground cost are unbounded, as it is the case in the landmark example of the quadratic cost with Gaussian marginals. This article aims at showing that semiconcavity bounds for entropic and Sinkhorn potentials play a key role in answering some of these questions. We now provide the reader with some background on EOT, and then proceed to describe our main contributions.
1.1. Background on EOT and Sinkhorn algorithm
The entropic optimal transport problem is a regularized version of the Monge-Kantorovich problem
(MK) $\inf_{\pi\in\Pi(\rho,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c\,\mathrm{d}\pi\,.$
In the case when $\mathcal{X}=\mathcal{Y}$ is endowed with a distance ${\sf d}$ and $c(x,y)={\sf d}(x,y)^{2}$, the optimal value of (MK) coincides with the squared Wasserstein distance of order two, which we denote by $\mathbf{W}_{2}^{2}(\mu,\nu)$. It is known [47, Thm 4.2] that under mild integrability conditions on the cost $c$, there exist two functions $\varphi^{\nu}_{\varepsilon}:\mathcal{X}\rightarrow\overline{\mathbb{R}},\psi^{\nu}_{\varepsilon}:\mathcal{Y}\rightarrow\overline{\mathbb{R}}$ with $\overline{\mathbb{R}}=\mathbb{R}\cup\{-\infty\}$, called entropic potentials, such that the unique optimal plan $\pi_{\varepsilon}^{\rho\nu}$ for (EOT) admits the Radon–Nikodým density
[…]
For any measurable $\varphi:\mathcal{X}\rightarrow\overline{\mathbb{R}}$ and $\psi:\mathcal{Y}\rightarrow\overline{\mathbb{R}}$ we set
(1.2) $\Phi^{\rho}_{0}(\varphi)(y)$ $:=-\varepsilon\log\int_{\mathcal{X}}\exp\biggl{(}-\frac{c(x,y)+\varphi(x)}{\varepsilon}\biggr{)}\rho(\mathrm{d}x)\,,$
$\Psi^{\nu}_{0}(\psi)(x)$ $:=-\varepsilon\log\int_{\mathcal{Y}}\exp\biggl{(}-\frac{c(x,y)+\psi(y)}{\varepsilon}\biggr{)}\nu(\mathrm{d}y)\,.$
Imposing that a probability measure of the form (1.1) belongs to $\Pi(\rho,\nu)$ yields the following non-linear system, also known as Schrödinger system,
[…]
Note that a priori the identities (1.3) are valid only $\rho$-almost surely and $\nu$-almost surely respectively. However, since $\Psi_{0}^{\nu}$ and $\Phi^{\rho}_{0}$ are well-defined even outside the supports, we obtain extensions of $\varphi^{\nu}_{\varepsilon}$, $\psi^{\nu}_{\varepsilon}$ taking values in ${}\mathbb{R}\cup\{-\infty\}$ such that (1.3) holds everywhere on $\mathcal{X}\times\mathcal{Y}$. Sinkhorn’s algorithm solves (1.3) as a fixed point problem, i.e., it constructs two sequences of potentials $(\varphi^{n}_{\varepsilon},\psi^{n}_{\varepsilon})$ defined through the iterations
(1.4) $\begin{cases}\varphi^{n+1}_{\varepsilon}&=-\Psi_{0}^{\nu}(\psi^{n}_{\varepsilon})\,,\\ \psi^{n+1}_{\varepsilon}&=-\Phi^{\rho}_{0}(\varphi^{n+1}_{\varepsilon})\,.\end{cases}$
Typically, the initialization is $\varphi^{0}_{\varepsilon}=0$, but other choices are possible.
1.2. Entropic stability for optimal plans
In this article, we consider a broad setting in which we require the target space $(\mathcal{Y},\mathfrak{g})$ to be a (possibly unbounded) connected Riemannian manifold without boundary endowed with its Riemannian metric $\mathfrak{g}$ whose associated Riemannian distance we denote ${\sf d}$. To fix ideas, one can simply take $\mathcal{Y}=\mathbb{R}^{d}$ equipped with the standard Euclidean metric. We will often consider this setting when applying the abstract results to concrete examples in Section 1.4. To state our main results, we need to introduce a notion of $\Lambda$-semiconcavity by saying that a differentiable function $f:(\mathcal{Y},\mathfrak{g})\rightarrow\mathbb{R}$ is $\Lambda$-semiconcave if for all $z,y\in\mathcal{Y}$ and for any geodesic $(\gamma_{t})_{t\in[0,1]}$ such that $\gamma_{0}=y$ and $\gamma_{1}=z$ we have
(1.5) $f(z)-f(y)\leq\langle\nabla f(y),\,\dot{\gamma_{0}}\rangle_{\mathfrak{g}}+\frac{\Lambda}{2}\,{\sf d}(z,y)^{2}\,.$
In the Euclidean setting, where geodesics are straight lines, this is equivalent to requiring that for all $z,y\in\mathcal{Y}$ it holds
$f(z)-f(y)\leq\langle\nabla f(y),\,z-y\rangle+\frac{\Lambda}{2}\,{\sf d}(z,y)^{2}\,.$
The first contribution of this article is an abstract $\mathrm{KL}$-stability bound for optimal plans associated with different marginal distributions.
Theorem 1.1 (KL stability of optimal plans).
Let $\pi_{\varepsilon}^{\rho\nu},\,\pi_{\varepsilon}^{\rho\mu}$ denote the unique optimizers in (EOT) for the set of marginals $(\rho,\nu)$ and $(\rho,\mu)$. If there exists $\Lambda>0$ such that
$y\mapsto c(x,y)+\psi^{\nu}_{\varepsilon}(y)$
is $\Lambda$-semiconcave uniformly in $x\in\operatorname{supp}(\rho)$, then
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\mu}|\pi_{\varepsilon}^{\rho\nu})\leq\mathrm{KL}(\mu|\nu)+\frac{\Lambda}{2{\varepsilon}}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,.$
The key step in the proof of Theorem 1.1 is the observation, made at Lemma 2.2, that the semiconcavity assumption implies the following $\mathrm{KL}$-bound between conditional distributions
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}(\cdot|y)|\pi_{\varepsilon}^{\rho\nu}(\cdot|z))\leq\frac{\Lambda}{2\varepsilon}\,{\sf d}^{2}(y,z)\,.$
The desired conclusion is then obtained using the gluing method, which is often employed in works on the stability of both entropic and classical optimal transport, see [29, 26, 44] for example.
Literature review: stability
Many recent works study the stability for entropic optimal transport with respect to variations in the marginal inputs. In [10], we obtain entropic stability bounds in a negative Sobolev norm, for a general class of problems with costs induced by diffusions on Riemannian manifolds with Ricci curvature bounded from below (which includes the quadratic cost on $\mathbb{R}^{d}$). The article [29] obtains a quantitative Hölder estimate between the Wasserstein distance of optimal plans and that of their marginals. This result applies to the quadratic cost, and to more general costs. Among the assumptions required, there is a transport-entropy inequality for marginals, which relates to Talagrand’s inequality, see TI($\tau$) below. A more qualitative stability is proven under mild hypothesis in [34]. On the dual side, Carlier and Laborde show in [17] $\mathrm{L}^{\infty}$-Lipschitz bounds for potentials for multimarginal problems in a bounded setting. The authors of [23] succeed in controlling the $\mathrm{L}^{\infty}$-norm of the difference of entropic potentials with the Wasserstein distance of order one between marginals using an approach based on Hilbert’s metric. Among their assumptions, there is boundedness of the cost function and compactness of the marginals’ supports. Subsequently, [11] bounds the difference of potentials with the Wasserstein distance of order two of the respective marginals. The norm used to express these bounds depends on the smoothness and boundedness of the cost. For example, the authors show an $\mathrm{L}^{\infty}$-bound between the gradient of entropic potentials provided the cost is bounded with two bounded derivatives. For the quadratic cost, gradients of entropic potentials provide with good approximations of optimal transport maps (see [36, 46] in unbounded settings and [51] in semidiscrete settings), thus justifying the growing interest around the stability of EOT. [26] establishes Lipschitz bounds between the $\mathrm{L}^{2}$-norm of the difference of the gradients of entropic potentials and the Wasserstein distance of the marginals. Here, the dependence of the Lipschitz constant on the regularization parameter is polynomial, thus improving on earlier results, and marginals may have unbounded support. All these results are obtained using a functional inequality for tilt-stable probability measures, see [14] and [2, Lemma 3.21]. Their main assumption is that both entropic potentials have a bounded Hessian. For this reason, these stability results are the closest in spirit to ours and partially inspired our work. Nonetheless, there are some important differences: we only require a bound on the Hessian of one of the two potentials, and the scope of Theorem 1.1 is not restricted to the quadratic cost, but covers general semiconcave costs. Finally, we mention the works [32, 44] since, even if they study the stability problem for unregularized optimal transport, they are somewhat connected to this work. In [32, Prop. 8] the authors establish a link between the optimality gap and the distance from the optimal plan, provided that the optimal transport map is generated by a $c$-concave potential. In [44] covariance inequalities are employed to bound the distance between transport maps. As we shall see in the sequel of this article (and as already understood in [20, 19]), these inequalities are valid tools for estimating the concavity properties of entropic potentials.
1.3. Exponential convergence of Sinkhorn’s algorithm
Sinkhorn’s algorithm admits a primal interpretation as the iterated Bregman projection algorithm for the relative entropy functional, see [3]. To explain this, we introduce two sequences of plans, called Sinkhorn plans, as follows:
(1.6) $\frac{\mathrm{d}\pi^{n,n}}{\mathrm{d}(\rho\otimes\nu)}=\exp\Big{(}-\frac{c+\varphi^{n}_{\varepsilon}\oplus\psi^{n}_{\varepsilon}}{\varepsilon}\Big{)},\quad\frac{\mathrm{d}\pi^{n+1,n}}{\mathrm{d}(\rho\otimes\nu)}=\exp\Big{(}-\frac{c+\varphi^{n+1}_{\varepsilon}\oplus\psi^{n}_{\varepsilon}}{\varepsilon}\Big{)}\,.$
In the above, and in the rest of the paper, for given potentials $\varphi:\mathcal{X}\rightarrow\mathbb{R}$, $\psi:\mathcal{Y}\rightarrow\mathbb{R}$ we set $\varphi\oplus\psi(x,y)=\varphi(x)+\psi(y)$. Then, $\pi^{n+1,n}$ may be viewed as the entropic projection of $\pi^{n,n}$ in the sense of Csiszár [21] on the set of plans having first marginal $\rho$. Likewise, $\pi^{n+1,n+1}$ is the entropic projection of $\pi^{n+1,n}$ on the set of plans having second marginal $\nu$. The second marginal $\nu^{n+1,n}$ of $\pi^{n+1,n}$ and the first marginal $\rho^{n,n}$ of $\pi^{n,n}$ are then given by
(1.7) $\mathrm{d}\nu^{n+1,n}\coloneqq e^{-\nicefrac{{(\psi^{n}_{\varepsilon}-\psi^{n+1}_{\varepsilon})}}{{\varepsilon}}}\,\mathrm{d}\nu\,,\quad\mathrm{d}\rho^{n,n}\coloneqq e^{-\nicefrac{{(\varphi^{n}_{\varepsilon}-\varphi^{n+1}_{\varepsilon})}}{{\varepsilon}}}\,\mathrm{d}\rho\,.$
The second general result of this article is a sufficient condition for the exponential convergence of Sinkhorn plans to the optimal plan $\pi_{\varepsilon}^{\rho\nu}$. To state it, let us recall that a probability measure $\nu\in\mathcal{P}(\mathcal{Y})$ is said to satisfy a Talagrand’s inequality with constant $\tau$, TI($\tau$) for short, if
(TI($\tau$)) $\mathbf{W}_{2}^{2}(\mu,\nu)\leq 2\tau\,\mathrm{KL}(\mu|\nu)\,,\quad\forall\mu\in\mathcal{P}(\mathcal{Y})\,.$
Theorem 1.2.
Assume that there exist $\Lambda\in(0,+\infty)$ and $N\geq 2$ such that
$y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$
is $\Lambda$-semiconcave uniformly in $x\in\operatorname{supp}(\rho)$ and $n\geq N$.
  1. (i)
    If there exists $\tau\in(0,+\infty)$ such that $\nu$ satisfies TI($\tau$), then for all $n\geq N-1$
    $\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n+1})\leq\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})\leq\bigg{(}1-\frac{\min\{\tau\Lambda,\,\varepsilon\}}{\min\{\tau\Lambda,\,\varepsilon\}+\tau\Lambda}\bigg{)}^{(n-N+2)}\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{0,0})\,.$
  2. (ii)
    If there exists $\tau\in(0,+\infty)$ such that $\nu^{n,n-1}$ satisfies TI($\tau$) for all $n\geq N$, then
    $\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n+1})\leq\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})\leq\bigg{(}1-\frac{\varepsilon}{\varepsilon+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{0,0})$
    holds for all $n\geq N$.
1.4. Examples of application
We now discuss several applications of Theorem 1.2. We emphasize that this result provides with a general sufficient condition for the exponential convergence of Sinkhorn’s algorithm: the examples we present here do not represent its full range of applicability, which deserves to be further explored.
  • If we write that “Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon\leq\varepsilon_{m}$ (explicitly given) the rate is $1-\kappa(\varepsilon)$ we always mean that for any $\varepsilon>0$ there is $\kappa^{\prime}\in(0,1)$ such that for all $n\geq 1$ we have
    (1.8) $\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n+1})\leq(1-\kappa^{\prime})^{n}\,\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{0,0})$
    and that if $\varepsilon\leq\varepsilon_{m}$, then $\kappa^{\prime}$ can be taken to be equal to $\kappa(\varepsilon)$. The requirement that $\varepsilon$ is small enough is just to have $\min\{\tau\Lambda,\varepsilon\}=\varepsilon$ when applying Theorem 1.2. In all examples, this hypothesis could be dropped, and the convergence rate is always explicit even for large $\varepsilon$, but we prefer to keep it for readability.
  • If we write that “Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon\leq\varepsilon_{m}$ (explicitly given) the asymptotic rate is $1-\kappa(\varepsilon)$ we mean that for any $\varepsilon>0$ (1.8) holds for some $\kappa^{\prime}\in(0,1)$ and all $n\geq 1$. Moreover, if $\varepsilon\leq\varepsilon_{m}$, for all $\delta>0$ there exists $N_{\delta}$ such that (1.8) holds for $n\geq N_{\delta}$ and $\kappa^{\prime}=\kappa(\varepsilon)-\delta$.
We start our gallery of examples with the case of (weakly) log-concave marginals. To state the next proposition, we introduce the notation $\|\Sigma\|_{2}$ to denote the operator norm of the matrix $\Sigma$.
[…]
Let $\mathcal{X},\mathcal{Y}\subseteq\mathbb{R}^{d}$ be two open (connected, possibly unbounded) domains of $\mathbb{R}^{d}$ endowed with the Euclidean metric and assume that $c(x,y)\coloneqq\langle x-y,\,\Sigma(x-y)\rangle/2$ for some positive definite symmetric matrix $\Sigma$. Further assume that $\nu$ satisfies TI($\tau$).
  1. (i)
    If $\rho(\mathrm{d}x)=e^{-U_{\rho}(x)}\mathrm{d}x$, $\nu(\mathrm{d}y)=e^{-U_{\nu}(y)}\mathrm{d}y$ and there exist $\alpha_{\rho},\beta_{\nu}\in(0,+\infty)$ such that
    (1.9) $\nabla^{2}U_{\rho}(x)\succeq\alpha_{\rho},\quad\nabla^{2}U_{\nu}(y)\preceq\beta_{\nu}\quad\forall x,y\in\mathbb{R}^{d}$
    in semidefinite order, then Sinkhorn’s algorithm converges exponentially fast. If we choose $\varphi^{0}_{\varepsilon}$ such that $\nabla^{2}\varphi^{0}_{\varepsilon}\succeq(\sqrt{\nicefrac{{\alpha_{\rho}}}{{\beta_{\nu}}}}-1)\Sigma$ [^1] In fact, the result holds under the weaker assumption $\nabla^{2}\varphi^{\varepsilon}_{0}\succeq-\Sigma-\varepsilon\frac{\alpha_{\rho}}{2}+\Big{(}\frac{\varepsilon^{2}\alpha^{2}_{\rho}}{4}+\frac{\alpha_{\rho}}{\beta_{\nu}}\Sigma^{2}\Big{)}^{1/2},$ but we prefer to keep the stronger condition for better readability. , then for $\varepsilon\leq\tau\,\sqrt{\nicefrac{{\beta_{\nu}}}{{\alpha_{\rho}}}}\,\left\|\Sigma\right\|_{2}$ the rate is
    (1.10) $1-\frac{\varepsilon}{\varepsilon+\tau\|\Sigma\|_{2}(\nicefrac{{\beta_{\nu}}}{{\alpha_{\rho}}})^{\nicefrac{{1}}{{2}}}}\,.$
    Otherwise, for any general initialization $\varphi^{0}_{\varepsilon}$ the expression appearing in (1.10) represents the asymptotic rate.
  2. (ii)
    If $\rho(\mathrm{d}x)=e^{-U_{\rho}(x)}\mathrm{d}x$ and there exists $\alpha_{\rho}\in(0,+\infty)$ such that
    $\nabla^{2}U_{\rho}(x)\succeq\alpha_{\rho}\quad\forall x\in\mathbb{R}^{d}$
    in semidefinite order, then Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon\leq\|\Sigma\|_{2}\,\sqrt{\nicefrac{{\tau}}{{\alpha_{\rho}}}}$ the rate is
    (1.11) $1-\frac{\varepsilon^{2}\,\alpha_{\rho}}{\varepsilon^{2}\,\alpha_{\rho}+\tau\,\|\Sigma\|_{2}^{2}}\,.$
  3. (iii)
    In the quadratic cost case $\Sigma=\operatorname{Id}$, if $\rho$ is weakly-log-concave of parameters $\alpha_{\rho}$ and $L$, then Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon\leq\alpha_{\rho}^{-1}\sqrt{\tau(\alpha_{\rho}+L)}$ the rate is
    $1-\frac{\varepsilon^{2}\alpha_{\rho}^{2}}{\varepsilon^{2}\alpha_{\rho}^{2}+\tau(\alpha_{\rho}+L)}\,.$
Following [12] we say that $\rho$ is weakly log-concave of parameters $\alpha_{\rho}>0$ and $L\geq 0$ if $\rho(\mathrm{d}x)=e^{-U_{\rho}(x)}\mathrm{d}x$, $U_{\rho}$ is continuously differentiable and
(1.12) $\langle\nabla U_{\rho}(\hat{x})-\nabla U_{\rho}(x),\hat{x}-x\rangle\geq\alpha_{\rho}|\hat{x}-x|^{2}-|\hat{x}-x|f_{L}(|\hat{x}-x|)\quad\forall x,\hat{x}\in\mathbb{R}^{d},$
where for any $L\geq 0$, $r>0$ we define $f_{L}(r)=2L^{1/2}\tanh\big{(}(L^{1/2}r)/2\big{)}$. If $L=0$, $\rho$ is $\alpha_{\rho}$-log-concave. If $L>0,$ then the Hessian of $U_{\rho}$ is lower bounded by $-L$ and $U_{\rho}$ behaves almost like an $\alpha_{\rho}$-convex potential for distant points. Weakly log-concave distributions form a rich class of probability measures that may be viewed as perturbations of log-concave distributions. For example, if $U_{\rho}$ is a double-well potential of the form
$U_{\rho}(x)=|x|^{4}-M|x|^{2}+C\,,$
for some $M>0$, $C\in\mathbb{R}$, then $\rho$ is weakly log-concave. More generally, if $U_{\rho}=V_{\rho}+W_{\rho}$ with $V_{\rho}$ strongly convex and $W_{\rho}$ Lipschitz with Lipschitz derivative, then $\rho$ is weakly log-concave.
The previous result applies to the fundamental example of the Euclidean quadratic cost $|x-y|^{2}$ as well as other costs considered in practical applications like the subspace elastic costs
(1.13) $c(x,y)=\frac{|x-y|^{2}}{2}+\frac{\gamma}{2}\,|A^{\perp}(x-y)|^{2}\,,$
where $A$ is a fixed $p$-rank matrix and $A^{\perp}$ its orthogonal projection $A^{\perp}=\operatorname{Id}-A^{\intercal}(AA^{\intercal})^{-1}A$ (if $A$ satisfies $AA^{\intercal}=\operatorname{Id}$, then $A^{\perp}=\operatorname{Id}-A^{\intercal}A$). More precisely, Proposition 1.4 applies to $\Sigma=\operatorname{Id}+\gamma\,A^{\perp}$ having norm equal to $\|\Sigma\|_{2}=1+\gamma\|A^{\perp}\|_{2}$. These costs have recently been considered in [16, 40] where the authors notice that they tend to promote sparsity for the corresponding optimal transport map. Let us also remark that the convergence rate obtained here for marginals satisfying (1.9) is tight in $\varepsilon$. Indeed, [13, Theorem 1.3] shows that the convergence rate is always bounded from below from $1-\varepsilon$ when considering Gaussian marginals (which satisfy (1.9)).
If we consider a marginal $\rho$ satisfying a light-tail condition, then we can address exponential convergence of more general costs. Here and below, for a function $u:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ we write $\nabla_{1}u(x,y)$, $\nabla_{1}^{2}u(x,y)$ for the gradient and the Hessian with respect to the first component and similarly for $\nabla_{2}u(x,y)$ and $\nabla_{2}^{2}u(x,y)$.
[…]
Let $\mathcal{X},\mathcal{Y}\subseteq\mathbb{R}^{d}$ be two open (connected, possibly unbounded) domains of $\mathbb{R}^{d}$ endowed with the Euclidean metric and assume that $\rho$ has light tails in the following sense: there exist $C,\delta>0$, $R\geq 0$, $L\geq 0$ such that
(1.14) $\nabla^{2}U_{\rho}(x)\succeq\begin{cases}C\,|x|^{\delta}\quad&\text{ for }|x|>R\,,\\ -L\quad&\text{ for }|x|\leq R\,.\end{cases}$
Then we distinguish two cases.
  1. (i)
    If $c(\cdot,\cdot)$ satisfies
    (1.15) $\nabla_{1}^{2}c(x,y),\nabla^{2}_{2}c(x,y)\succeq h(c)\quad\text{ and }\quad\left\|\nabla_{1}\nabla_{2}c(x,y)\right\|_{2}\vee\left\|\nabla_{1}^{2}c(x,y)\right\|_{2}\leq H(c)\,$
    uniformly in $x,y\in\mathbb{R}^{d}$, for some constants $h(c),\,H(c)\in\mathbb{R}$, then Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon$ small enough (as in (2.27)) the rate is
    $1-\frac{\varepsilon^{2+\nicefrac{{2}}{{\delta}}}}{\varepsilon^{2+\nicefrac{{2}}{{\delta}}}+\varepsilon^{1+\nicefrac{{2}}{{\delta}}}\tau\delta_{H}+2\tau H(c)^{2}\,(\varepsilon^{\nicefrac{{2}}{{\delta}}}+(L+2)\,C^{-\nicefrac{{2}}{{\delta}}}(\varepsilon+\delta_{H})^{\nicefrac{{2}}{{\delta}}})}$
    where $\delta_{H}\coloneqq H(c)-h(c)$.
  2. (ii)
    When $c(x,y)\coloneqq\langle x-y,\,\Sigma(x-y)\rangle/2$ is the (anisotropic) quadratic cost, Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon$ small enough (as in (2.28)) the rate is
    $1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\,\tau\,\|\Sigma\|_{2}^{2}\,\Bigl{(}1+(L+2)\bigl{[}R^{2}\vee{C^{-\nicefrac{{2}}{{\delta}}}}\bigr{]}\Bigr{)}}\,.$
The class of probability measures satisfying the light-tails condition (1.14) includes for instance the potentials with tails lighter than Gaussian, e.g. $U_{\rho}(x)=C_{0}(1+|x|^{2+\delta})$ for which $C=C_{0}(1+\delta)(2+\delta)$ and $L=R=0$. The convergence stated above applies to a general class of costs including $p$-costs $c(x,y)=(1+|x-y|^{2})^{\nicefrac{{p}}{{2}}}-1$, with $p\in(1,2)$, and the STVS [^2] Soft-Thresholding operator with Vanishing Shrinkage. elastic cost (as proposed in [55], see also [40]), that is the cost
$c(x,y)=\frac{|x-y|^{2}}{2}+\gamma^{2}\,\sum_{i=0}^{d}\biggl{(}\mbox{asinh}\biggl{(}\frac{|x_{i}-y_{i}|}{2\gamma}\biggr{)}+\frac{1}{2}-\frac{1}{2}e^{-2\,\mbox{asinh}\left(\frac{|x_{i}-y_{i}|}{2\gamma}\right)}\biggr{)}\,,$
for which for any $j\in\{1,2\}$ we have
$\frac{1}{2}\preceq\nabla_{j}^{2}c(x,y)=\operatorname{Id}+\frac{1}{2}\mathrm{diag}\biggl{(}\frac{|x_{i}-y_{i}|}{\sqrt{(x_{i}-y_{i})^{2}+\gamma^{2}}}\biggr{)}-\frac{1}{2}\preceq 1\,,$
and similarly $\left\|\nabla_{1}\nabla_{2}c(x,y)\right\|_{2}\leq\nicefrac{{3}}{{2}}$, and thus we get a rate independently of the regularizing parameter $\gamma$. To the best of our knowledge, the only exponential convergence results that apply under the assumptions of Proposition 1.4 are in [28]. While the assumptions made therein are weaker, we are able to improve on the convergence rate by showing that it is polynomial in $\varepsilon$.
For Lipschitz costs, we deduce the following convergence result.
[…]
Let $\mathcal{X},\mathcal{Y}\subseteq\mathbb{R}^{d}$ be two open (connected, possibly unbounded) domains of $\mathbb{R}^{d}$ endowed with the Euclidean metric and assume that there are two constants $h(c),\,H(c)\in\mathbb{R}$ such that
(1.16) $h(c)\preceq\nabla^{2}_{2}c(x,y)\preceq H(c)\quad\text{ and }\quad\nabla^{2}_{1}c(x,y)\preceq H(c)\,\quad\forall x,y\in\mathbb{R}^{d}.$
We distinguish two cases.
  1. (i)
    Assume that the cost is Lipschitz in $y$, uniformly in $x$, that is,
    (1.17) $\sup_{x\in\mathbb{R}^{d}}|c(x,y)-c(x,\hat{y})|\leq\mathrm{Lip}_{\infty,2}(c)\,|y-\hat{y}|\,\quad\forall y,\hat{y}\in\mathbb{R}^{d}.$
    Then, Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon$ small enough (as in (2.29)) the rate is
    $1-\frac{\varepsilon^{2}}{\varepsilon^{2}+\tau\varepsilon(H(c)-h(c))+\tau\mathrm{Lip}^{2}_{\infty,2}(c)}\,.$
  2. (ii)
    Assume that the cost is Lipschitz in $x$, uniformly in $y$, that is,
    (1.18) $\sup_{y\in\mathbb{R}^{d}}|c(x,y)-c(\hat{x},y)|\leq\mathrm{Lip}_{1,\infty}(c)\,|x-\hat{x}|\ \quad\forall x,\hat{x}\in\mathbb{R}^{d}\,.$
    Further assume that $\left\|\nabla_{12}^{2}c(x,y)\right\|_{2}\leq H(c)$ and that $\rho$ satisfies the logarithmic Sobolev inequality LSI($C_{\rho}$). Then, Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon$ small enough (as in (2.30)) the rate is
    $1-\frac{\varepsilon^{4}}{\varepsilon^{4}+\tau\varepsilon^{3}(H(c)-h(c))+2\tau H(c)^{2}C_{\rho}(\varepsilon^{2}+4C_{\rho}\mathrm{Lip}_{1,\infty}^{2})}\,.$
We recall here that $\rho$ satisfies a logarithmic Sobolev inequality with constant $C_{\rho}$, LSI($C_{\rho}$) for short, if for any probability measure $\mathfrak{p}\in\mathcal{P}(\mathcal{X})$ we have
(LSI($C_{\rho}$)) $\mathrm{KL}(\mathfrak{p}|\rho)\leq\frac{C_{\rho}}{2}\int\left|\nabla\log\frac{\mathrm{d}\mathfrak{p}}{\mathrm{d}\rho}\right|^{2}\mathrm{d}\mathfrak{p}\,.$
Many probability measures satisfy LSI, for instance that is the case if $U_{\rho}$ is $C_{\rho}^{-1}$-convex, or weakly asymptotically convex (see [5] for more general measures). As a direct consequence of Proposition 1.4, when considering symmetric costs that are Lipschitz, one obtains the exponential convergence of Sinkhorn’s algorithm with a rate that is quadratic in $\varepsilon$, without any assumption other than TI($\tau$) for $\nu$. We are not aware of any exponential convergence result valid under the assumptions of (i). On the contrary, if we consider the setting of (ii), then the results of [28] apply, since both TI($\tau$) and LSI($C_{\rho}$) imply Gaussian tails. However, we improve the dependence of the rate of convergence in $\varepsilon$, which we show to be polynomial.
Theorem 1.2 can also be employed in the compact setting, leading to the following.
[…]
Assume that $\rho$ has compact support and that $\nu$ satisfies TI($\tau$).
  1. (i)
    If $\nabla_{2}^{2}c(x,y)=\Sigma$, then Sinkhorn’s algorithm converges and for $\varepsilon\leq\sqrt{\tau}\,\left\|g\right\|_{\mathrm{L}^{\infty}(\rho)}$ the rate is
    (1.19) $1-\frac{\varepsilon^{2}}{\varepsilon^{2}+\tau\|g\|_{\mathrm{L}^{\infty}(\rho)}^{2}}\,,$
    where $g(x)\coloneqq\nabla_{2}c(x,y)-\Sigma y$.
  2. (ii)
    If $\nu$ is compactly supported, then Sinkhorn’s algorithm converges and for $\varepsilon$ small enough (as in (2.31)) the rate is
    $1-\frac{\varepsilon^{2}}{\varepsilon^{2}+\tau\varepsilon\,(H(c)-h(c))+\tau\left\|\nabla_{2}c\right\|_{\mathrm{L}^{\infty}(\rho\times\nu)}^{2}}\,,$
    where $h(c),\,H(c)$ are defined such that $h(c)\preceq\nabla_{2}^{2}c(x,y)\preceq H(c)$ on $\operatorname{supp}(\rho)\times\operatorname{supp}(\nu)$.
As a straightforward consequence, in the case $c(x,y)=\nicefrac{{|x-y|^{2}}}{{2}}$, if $\operatorname{supp}(\rho)\subseteq B_{R}(0)$, then the convergence rate is $1-\nicefrac{{\varepsilon^{2}}}{{\varepsilon^{2}+\tau R^{2}}}$. The only exponential convergence results with polynomial rates in a setting comparable to that of Proposition 1.4-(i) are those obtained in [13, Theorem 1.2]. The authors impose that one the two marginals has a density bounded above and below and that its support is compact and convex. This hypothesis is stronger than assuming that $\nu$ satisfies a Talagrand’s inequality. Indeed, it implies LSI and hence TI, cf. [6, Proposition 5.3] and discussion therein, and [5, Proposition 5.1.6]. About convergence rates, in [13] the authors obtain a rate which behaves like $\varepsilon^{2}$ for small $\varepsilon$, which is the same behavior of (1.19). Moreover, they also show that the asymptotic rate is of order $\varepsilon$. However, this result is not obviously comparable to the findings of Proposition 1.4, as the number of iterations after which this rate is shown to hold diverges as $\varepsilon\rightarrow 0$. The findings of Proposition 1.4-(ii) are to be compared with those of [13, Theorem 3.1-(A1)]. As before, we can drop the assumption that the support is convex and the same considerations made in the previous case regarding convergence rates hold.
Lastly, we discuss the convergence of Sinkhorn’s algorithm on Riemannian manifolds, starting from the $d$-dimensional Riemannian sphere $\mathbb{S}^{d}\subseteq\mathbb{R}^{d+1}$ endowed with the angular metric and corresponding angular distance between two points, that is defined for any $x,\,y\in\mathbb{S}^{d}$ as
${\sf d}(x,y)=\arccos(\langle x,y\rangle)\,.$
On ${\mathbb{S}^{d}}$ we will consider the EOT problem associated to two possible cost functions
$c(x,y)=1-\cos({\sf d}(x,y))=1-\langle x,y\rangle\,,\quad\text{ and }\quad c_{\delta}(x,y)=\arccos(\delta\langle x,y\rangle)^{2}$
with $\delta\in(0,1)$ fixed. Notice that for $\delta=1$ we have $c_{\delta}(x,y)={\sf d}(x,y)^{2}$, so that $c_{\delta}$ can be seen as a smoothed version of the cost ${\sf d}^{2}$ which avoids the singularity the Hessian of ${\sf d}^{2}$ faces when considering antipodal points $(x,-x)$. For this class of problems we can prove the following.
[…]
Assume that $\nu$ satisfies TI($\tau$). We distinguish two cases.
  1. (i)
    When considering the regular cost $c(x,y)=1-\langle x,y\rangle$, Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon\leq 1+\sqrt{1+\tau^{2}}$ the exponential convergence rate equals
    $1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\tau\varepsilon+\tau}\,.$
  2. (ii)
    When the cost considered is $c_{\delta}(x,y)$, Sinkhorn’s algorithm converges exponentially fast and for $\varepsilon$ small enough (as in (2.35)) the rate of convergence is
    $1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\tau\varepsilon\left(\delta^{2}+\frac{4\pi}{\sqrt{1-\delta^{2}}}\right)+\frac{8\pi^{2}\tau}{1-\delta^{2}}}\,.$
As a final example, we show that our result also applies to the classical Schrödinger problem [54, 43] on any compact smooth Riemannian manifold $(M,g)$. This problem is usually written as the entropy minimization
(1.20) $\inf_{\pi\in\Pi(\rho,\nu)}\mathrm{KL}(\pi|R_{0\varepsilon})$
where $R_{0\varepsilon}(\mathrm{d}x\mathrm{d}y)=\mathsf{p}_{\varepsilon}(x,y){\rm vol}(\mathrm{d}x){\rm vol}(\mathrm{d}y)$, $\mathsf{p}_{\varepsilon}(x,y)$ being the heat kernel (namely the unique solution at time $\varepsilon$ of the ‘heat equation’ $\partial_{t}u=\frac{1}{2}\Delta u$ for the initial condition $u(0,x)=\delta_{x}$) and ${\rm vol}$ the volume form. As explained for instance in [10, Section 3], (1.20) can be recast as an EOT problem with cost function
$c_{\varepsilon}(x,y)=-\varepsilon\log\mathsf{p}_{\varepsilon}(x,y)\,.$
Since upper and lower bounds on the sectional curvatures of $M$ allow to control the Hessian of $\log\mathsf{p}_{t}$, we obtain the following
[…]
Let $M$ be a compact smooth Riemannian manifold and assume that $\nu$ satisfies TI($\tau$). Then for every $\varepsilon\in(0,1]$ Sinkhorn’s algorithm converges exponentially fast. Moreover, as soon as $\varepsilon$ is small enough (cf. (2.38)) the exponential convergence rate equals
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\tau C^{\prime}(\varepsilon+{\rm diam}(M)^{2})+\tau C(\varepsilon+\kappa^{-}\varepsilon(1+\varepsilon)+1)}$
where $\kappa\in\mathbb{R}$ denotes a lower bound for the Ricci curvature and $C,C^{\prime}$ are constants depending on the dimension, Ricci and sectional curvatures of $M$. Finally, if $M$ has non-negative Ricci curvature, then the exponential convergence of Sinkhorn’s algorithm holds for any regularizing parameter $\varepsilon>0$.
1.5. Literature review: Sinkhorn’s algorithm
Bounded costs.
Sinkhorn’s algorithm has a long history, going back at least to the works of Sinkhorn [56] and Sinkhorn and Knopp [57] in the discrete setting. Here, it is employed as an algorithm to construct a matrix with prescribed rows and columns sums. Other important early contributions include [31, 7]. In particular, [31] introduced Hilbert’s metric as a tool to prove exponential convergence. The realization [22] that EOT provides with a numerically more tractable version of the Monge-Kantorovich problem triggered an explosion of interest on the subject. For bounded costs, [15] obtained the first exponential convergence results in the continuous setting using the Hilbert metric approach and [25] establishes qualitative $\mathrm{L}^{p}$ convergence and regularity estimates for Sinkhorn iterates using an optimal transport approach; in particular these results apply to the multimarginal case as well. The exponential convergence of the algorithm in the multimarginal setting is a result of Carlier [9]. A more probabilistic viewpoint is introduced in [35], where contraction estimates are obtained by means of coupling arguments. The work [13] shows that exponential convergence takes place for semiconcave bounded costs, under various sets of hypotheses on the marginals. The main innovation of this work consists in showing that the exponential rate of convergence deteriorates polynomially in $\varepsilon$. This is in contrast with previous works that exhibited a rate of convergence that decays exponentially in $\varepsilon$.
Unbounded costs.
For unbounded costs and marginals, Rüschendorf [52] establishes qualitative convergence for Sinkhorn plans in relative entropy. These results are improved in [48] where the authors manage to show qualitative convergence on both the primal and dual side under mild assumptions on the cost and marginals. The work by Léger [42] provides with a insightful interpretation of Sinkhorn’s algorithm as a block-coordinate descent algorithm on the dual problem, see also [1, 41]. From this interpretation, it follows that the speed of convergence is at least $n^{-1}$ under minimal assumptions. Polynomial rates of convergence are also established in [29] as a consequence of the above mentioned stability results and subsequently improved in [33], exploiting a symmetrized version of the KL-divergence. It is only very recently that the first exponential convergence results for unbounded costs and marginals have appeared. To the best of our knowledge, the first article containing such results is [12], which studies the quadratic cost. The main result is that if the marginals are weakly log-concave and $\varepsilon$ is large enough, exponential convergence of $\nabla\varphi^{n}$ to $\nabla\varphi^{\nu}_{\varepsilon}$ takes place. This kind of convergence is particularly useful as $\nabla\varphi^{\nu}_{\varepsilon}$ approximates the Brenier map in the $\varepsilon\rightarrow 0$ limit, see [10, 51]. The proof follows a geometric approach that highlights that semiconvexity and semiconcavity bounds on Sinkhorn potentials can be leveraged to obtain exponential convergence. Subsequently, the article [28] succeeds in constructing versions of the Hilbert metric that are contractive for general unbounded costs. In contrast with [12], exponential convergence is shown for all values of $\varepsilon$. Roughly speaking, the main assumption is that there exist some $p>0$ such that $c(x,y)$ grows no faster than $|x|^{p}+|y|^{p}$ and the tails of both $\mu$ and $\nu$ decay faster than $\exp(-r^{p+\delta})$ for some $\delta>0$. When applied to the quadratic cost, this assumptions does not completely cover log-concave distributions and their perturbations, leaving out Gaussian marginals for example. Over the past few years, a number of relevant contributions focused on different asymptotic properties of Sinkhorn’s algorithm than the speed of convergence. It would be impossible to account for all of them here. Let us just mention [4] for the relation with Monge-Ampère equation, [24] for the construction of Wasserstein mirror gradient flows, and [53] for construction of a Transformer variant inspired by Sinkhorn’s algorithm.
1.5.1. A more precise comparison for the Euclidean quadratic cost
Here we compare the convergence rates obtained for the quadratic cost setting in this paper with the most recent exponential convergence rates deduced in [12, 28, 13]. The forthcoming discussion is summarized in Table 1.
Assumptions [12] [28] [13] Our results ($\nu\in$ TI($\tau$))
$\rho$ is $\alpha_{\rho}$-log-concave $\nu$ is $\beta_{\nu}$-log-convex $1-\Theta(\varepsilon)$ if $\mu,\nu$ are Gaussians $\varepsilon>\varepsilon_{0}$ if also $\nu$ log-concave $1-\Theta(\frac{\varepsilon}{\tau}\,\sqrt{\nicefrac{{\alpha_{\rho}}}{{\beta_{\nu}}}})$
$\rho$ is $\alpha_{\rho}$-log-concave $\varepsilon>\varepsilon_{0}$ if also $\nu$ weak log-concave $1-\Theta(\nicefrac{{\varepsilon}}{{R^{2}}})$ if $\operatorname{supp}(\rho),\operatorname{supp}(\nu)\subseteq B_{R}(0)$ and $\operatorname{supp}(\rho)$ is convex $1-\Theta(\varepsilon^{2}\,\nicefrac{{\alpha_{\rho}}}{{\tau}})$
$\rho$ is $(\alpha_{\rho},L)$-weak log-concave $\varepsilon>\varepsilon_{0}$ if also $\nu$ weak log-concave $1-\Theta(\nicefrac{{\varepsilon^{2}\alpha_{\rho}^{2}}}{{(\alpha_{\rho}+L)\tau}})$
$\rho$ with light tails $\varepsilon>\varepsilon_{0}$ if also $\nu$ has light tails $1-\Theta(e^{-\varepsilon^{-1}})$ if also $\nu$ has light tails weaker notion of light tails $1-\Theta(\nicefrac{{\varepsilon^{2}}}{{\tau(1+L\,[R^{2}\vee C^{-\nicefrac{{2}}{{\delta}}}])}})$
$\operatorname{supp}(\rho)\subseteq B_{R}(0)$ $1-\Theta(e^{-\varepsilon^{-1}})$ if also $\nu$ has light tails weaker notion of light tails $1-\Theta(\nicefrac{{\varepsilon^{2}}}{{R^{4}}})$ if also $\operatorname{supp}(\nu)\subseteq B_{R}(0)$, $\operatorname{supp}(\rho)$ is convex and $m\leq\log\rho(x)\leq M$ $1-\Theta(\nicefrac{{\varepsilon^{2}}}{{\tau R^{2}}})$
Table 1. Comparison of rates for the Euclidean setting with quadratic cost. We always assume $\nu$ to satisfy the transport inequality TI($\tau$). In the table $\Theta(f)$ is a function for which there are universal constants $c,C>0$ such that $cf\leq\Theta(f)\leq Cf$.
The main novelties that we introduce in the quadratic setting are the following.
  • We obtain for the first time exponential convergence for all $\varepsilon>0$ under the assumption that $\rho$ is $\alpha_{\rho}$-log-concave, $\nu$ is $\beta_{\nu}$-log-semiconvex and satisfies a Talagrand inequality. The convergence rate we obtain is $1-\Theta(\frac{\varepsilon}{\tau}\,\sqrt{\nicefrac{{\alpha_{\rho}}}{{\beta_{\nu}}}})$, which has a sharp dependence on $\varepsilon$ thanks to [13, Theorem 1.3]. The only result that covers this setup is obtained in [12] but it gives exponential convergence only for $\varepsilon$ large enough.
  • We obtain for the first time exponential convergence assuming that $\rho$ is $\alpha_{\rho}$-log-concave and $\nu$ satisfies TI($\tau$). The convergence rate is $1-\Theta(\varepsilon^{2}\,\nicefrac{{\alpha_{\rho}}}{{\tau}})$. There are two results we might compare with ours. In the unbounded setting [12] proves convergence for $\varepsilon>\varepsilon_{0}$ (with the noise threshold $\varepsilon_{0}$ being zero solely in the Gaussian setting). The second one is [13] where the authors manage to get a rate which in linear in $\varepsilon$ but have to assume that both $\rho$ and $\nu$ have compact support, which we do not. This rate has better dependence in $\varepsilon$ but it deteriorates with the size of the support. Our result shows that indeed stronger log-concavity of $\rho$ improves the convergence rate, addressing an issue raised in [13, Remark 3.3].
  • For weakly log-concave distributions we obtain for the first time exponential convergence for all $\varepsilon>0$. The only previously known result [12] requires $\varepsilon$ to be larger than a threshold value $\varepsilon_{0}$. Moreover, we show that the rate of convergence $1-\Theta(\nicefrac{{\varepsilon^{2}\alpha_{\rho}^{2}}}{{L\tau}})$ without having to assume neither compactness nor convexity of the marginals’ supports.
  • In the light-tails setting, we improve the rate of convergence, from an exponential dependence on $\varepsilon$ in [28], to the polynomial $1-\Theta(\nicefrac{{\varepsilon^{2}}}{{\tau(1+L\,[R^{2}\vee C^{-\nicefrac{{2}}{{\delta}}}])}})$. Moreover, we ask that only one of two marginals has tails lighter than Gaussian, whereas [28] requires this property to hold for both marginals. However, one should note that the result of [28] applies to more general costs and marginals. For example, the light-tail condition considered therein can be expressed as a condition on the growth $U_{\rho}$ rather than its Hessian. Moreover, the cost is not required to have a bounded Hessian.
  • When dealing with $\rho$ compactly supported, we show a rate $1-\Theta(\nicefrac{{\varepsilon^{2}}}{{\tau R^{2}}})$ under the assumption that both marginals have compact support and one of them satisfies TI($\tau$). This should be compared to [13, Theorem 3.1-(A1)]. There, the authors get the same dependence in $\varepsilon^{2}$ assuming that one marginal is compactly supported, and the other one has a uniformly upper- and lower-bounded density on its support, that is taken to be compact and convex. This last assumption implies LSI and is therefore stronger than asking that TI holds, cf. [6, Proposition 5.3], discussion therein and [5, Proposition 5.1.6]. Indeed, if LSI holds, TI also holds. As explained in the discussion following Proposition 1.4, the authors also prove that the asymptotic rate of convergence is of order $\varepsilon$.
The proof of all the above propositions is obtained bounding from above the Hessian of Sinkhorn potentials uniformly in $n$ and then invoking Theorem 1.2. To control Hessians, we leverage their representation in terms of conditional covariances (cf. (2.2) below). Then, in most cases we proceed to bound covariances by means of functional inequalities and perturbative arguments. In the case of log-concave and weakly log-concave marginals, we argue differently by showing that the map $\Psi^{\nu}_{0}(\cdot)$ preserves concavity, and that $\Phi^{\rho}_{0}(\cdot)$ preserves convexity and weak convexity. To do so, we rely on Prékopa–Leindler inequality following [30, 20] when assuming strong log-concavity, and on the more probabilistic constructions of [19] when assuming weak log-concavity.
2. proofs
2.1. Preliminaries
For the proof of Theorem 1.1, We shall need that the conditional distribution of $\pi_{\varepsilon}^{\rho\nu}$ with respect to the second component can be written as
(2.1) $\pi_{\varepsilon}^{\rho\nu}(\mathrm{d}x|y)$ $=\exp\biggl{(}-\frac{c(x,y)+\varphi^{\nu}_{\varepsilon}(x)+\psi^{\nu}_{\varepsilon}(y)}{\varepsilon}\biggr{)}\rho(\mathrm{d}x)\,.$
Note that in principle the conditional distribution is only defined $\nu$-a.e. Nonetheless, under the semiconcavity assumption of Theorem 1.1, one can see that $\psi_{\varepsilon}^{\nu}$ is locally bounded from below, and thus we can use the right hand side of (2.1) to extend the definition of the conditional measure to the whole $\mathcal{Y}$. The same considerations apply to Sinkhorn plans in all examples. That is to say
$\pi^{n,n}(\mathrm{d}x|y)$ $=\exp\biggl{(}-\frac{c(x,y)+\varphi^{n}_{\varepsilon}(x)+\psi^{n}_{\varepsilon}(y)}{\varepsilon}\biggr{)}\rho(\mathrm{d}x)\,,$
$\pi^{n,n}(\mathrm{d}y|x)$ $=\exp\biggl{(}-\frac{c(x,y)+\psi^{n}_{\varepsilon}(y)+\varphi^{n}_{\varepsilon}(x)}{\varepsilon}\biggr{)}\nu(\mathrm{d}y)\,.$
are versions of the conditional probability defined everywhere on $\mathcal{Y}$ and $\mathcal{X}$ respectively. We will use at several places the following identities (see [20, Lemma 1] and [18, Appendix A]) along Sinkhorn’s algorithm
(2.2) $\nabla\psi^{n}_{\varepsilon}(y)$ $=-\int_{\mathcal{X}}\nabla_{2}c(x,y)\,\pi^{n,n}(\mathrm{d}x|y)\,,$
$\nabla^{2}\psi^{n}_{\varepsilon}(y)$ $=-\int_{\mathcal{X}}\nabla_{2}^{2}c(x,y)\,\pi^{n,n}(\mathrm{d}x|y)+\varepsilon^{-1}\,\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\,.$
Similarly, whenever also $(\mathcal{X},\mathfrak{g})$ is a Riemannian manifold [^3] The identities (2.3) are not employed in the proofs of Theorem 1.1 and Theorem 1.2 where $\mathcal{X}$ can be taken to be Polish. These identities will be employed in some examples below, where we assume $\mathcal{X}$ having a Riemannian structure. we have
(2.3) $\nabla\varphi^{n+1}_{\varepsilon}(x)$ $=-\int_{\mathcal{Y}}\nabla_{1}c(x,y)\,\pi^{n,n}(\mathrm{d}y|x)\,,$
$\nabla^{2}\varphi^{n+1}_{\varepsilon}(x)$ $=-\int_{\mathcal{Y}}\nabla_{1}^{2}c(x,y)\,\pi^{n,n}(\mathrm{d}y|x)+\varepsilon^{-1}\,\mathrm{Cov}_{Y\sim\pi^{n,n}(\cdot|x)}(\nabla_{1}c(x,Y))\,.$
In order to guarantee the validity of the differentiation under the integral sign is enough starting Sinkhorn’s algorithm with $\varphi^{0}_{\varepsilon}$ smooth enough. These estimates actually hold true also for the limit entropic potentials in a weak sense. A proof of this, for our general setting, can be found in [18, Proposition 4.4 and Lemma 4.5] where the authors show that $\varphi^{\nu}_{\varepsilon}$ and $\psi^{\nu}_{\varepsilon}$ are $\mathcal{C}^{1}$ with Hölder continuous first derivative, the gradient identities hold true whereas the Hessian identities hold in a weak sense.
2.2. Proof of Theorem 1.1 and Theorem 1.2
In this section we provide the proofs of Theorem 1.1 and Theorem 1.2. The key estimate required for Theorem 1.1 is contained in Section 2.2, which bounds the relative entropy of the conditional distribution of the entropic plan at two different points by their distance squared.
[…]
Under the assumptions of Theorem 1.1, let $\pi_{\varepsilon}^{\rho\nu}$ be the entropic plan between $\rho$ and $\nu$ and let $\pi_{\varepsilon}^{\rho\nu}(\cdot|y)$ denote the conditional distribution (conditioned on the second variable being equal to $y$). Then for all $y,z\in\mathcal{Y}$ we have
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}(\cdot|y)|\pi_{\varepsilon}^{\rho\nu}(\cdot|z))\leq\frac{\Lambda}{2\varepsilon}\,{\sf d}^{2}(y,z)\,.$
Proof.
Let $y,z$ be given and define the function
(2.4) $g(x,\bar{y})=c(x,\bar{y})+\psi^{\nu}_{\varepsilon}(\bar{y})$
By assumption, $\bar{y}\mapsto g(x,\bar{y})$ is $\Lambda$-semiconcave uniformly in $x\in\operatorname{supp}(\rho)$. From the representation of conditional distributions (2.1) we immediately get
(2.5) $\varepsilon\,\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}(\cdot|y)|\pi_{\varepsilon}^{\rho\nu}(\cdot|z)))$ $=\int_{\mathcal{X}}g(x,z)-g(x,y)\,\pi_{\varepsilon}^{\rho\nu}(\mathrm{d}x|y)$
$\leq\left\langle\dot{\gamma}_{0},\,\int_{\mathcal{X}}\nabla_{2}g(x,y)\,\pi_{\varepsilon}^{\rho\nu}(\mathrm{d}x|y)\right\rangle_{\!\!\mathfrak{g}}+\frac{\Lambda}{2}\,{\sf d}^{2}(y,z)\,,$
where $(\gamma_{t})_{t\in[01]}$ is a geodesic from $y$ to $z$. Next, from the representation (2.2) of $\nabla\psi^{\nu}_{\varepsilon}(y)$ as conditional expectation we find
$\nabla_{2}g(x,y)=\nabla_{2}c(x,y)-\int_{\mathcal{X}}\nabla_{2}c(\bar{x},y)\,\pi_{\varepsilon}^{\rho\nu}(\mathrm{d}\bar{x}|y)\,,$
from which it follows that
$\int_{\mathcal{X}}\nabla_{2}g(x,y)\,\pi_{\varepsilon}^{\rho\nu}(\mathrm{d}x|y)=0\,.$
Using this identity in (2.5) gives the desired result. ∎
Proof of Theorem 1.1.
We assume without loss of generality that $\mathrm{KL}(\mu|\nu),\,\Lambda,\,\mathbf{W}_{2}^{2}(\mu,\nu)$ are all finite, otherwise there is nothing to prove, in particular $\mu$ is absolutely continuous with respect to $\nu$. First, we recall that $\pi_{\varepsilon}^{\rho\mu}$ can be seen as a Schrödinger optimal plan w.r.t. the reference measure $\pi_{\varepsilon}^{\rho\nu}$ for the Schrödinger problem, that is the (unique minimizer) for
(2.6) $\mathrm{KL}(\pi_{\varepsilon}^{\rho\mu}|\pi_{\varepsilon}^{\rho\nu})=\min_{\pi\in\Pi(\rho,\mu)}\mathrm{KL}(\pi|\pi_{\varepsilon}^{\rho\nu})\,.$
This directly follows from [47, Theorem 2.1.b] after noticing that
[…]
and hence also $\pi_{\varepsilon}^{\rho\nu}$-a.s. (since $\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\rho\otimes\nu)<\infty$). We now proceed to bound $\mathrm{KL}(\pi_{\varepsilon}^{\rho\mu}|\pi_{\varepsilon}^{\rho\nu})$ exhibiting a suitable admissible plan in (2.6). To this aim, fix a coupling $\tau\in\Pi(\mu,\nu)$ between $\mu$ and $\nu$ and let us consider the coupling
$\pi(\mathrm{d}x,\mathrm{d}y)\coloneqq\mu(\mathrm{d}y)\int_{\mathcal{Y}}\pi_{\varepsilon}^{\rho\nu}(\mathrm{d}x|z)\tau(\mathrm{d}z|y)\,.$
Notice that $\pi\in\Pi(\rho,\mu)$ and therefore from (2.6) it follows $\mathrm{KL}(\pi_{\varepsilon}^{\rho\mu}|\pi_{\varepsilon}^{\rho\nu})\leq\mathrm{KL}(\pi|\pi_{\varepsilon}^{\rho\nu})$. From the disintegration property of relative entropy (see for instance [43, Appendix A]) and from its convexity we may deduce that
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\mu}|\pi_{\varepsilon}^{\rho\nu})$ $\,\leq\mathrm{KL}(\pi|\pi_{\varepsilon}^{\rho\nu})\leq\mathrm{KL}(\mu|\nu)+\int_{\mathcal{Y}}\mathrm{KL}(\pi(\cdot|y)|\pi_{\varepsilon}^{\rho\nu}(\cdot|y))\mu(\mathrm{d}y)$
$\,\leq\mathrm{KL}(\mu|\nu)+\int_{\mathcal{Y}}\int_{\mathcal{Y}}\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}(\cdot|z)|\pi_{\varepsilon}^{\rho\nu}(\cdot|y))\tau(\mathrm{d}z|y)\mu(\mathrm{d}y)$
[…]
The desired conclusion follows by optimizing over $\tau\in\Pi(\mu,\nu)$. ∎
Proof of Theorem 1.2.
We start recalling the known identity
(2.7) $\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})-\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n,n-1})$ $=\frac{1}{\varepsilon}\int(\varphi^{n+1}_{\varepsilon}-\varphi^{n}_{\varepsilon})\oplus(\psi^{n}_{\varepsilon}-\psi^{n-1}_{\varepsilon})\,\mathrm{d}\pi_{\varepsilon}^{\rho\nu}$
=(1.7)${{=}}-(\mathrm{KL}(\rho|\rho^{n,n})+\mathrm{KL}(\nu|\nu^{n,n-1}))\,.$ ) .
Proof of (i). Let $n\geq N-1.$ The standing assumptions grant that we can apply Theorem 1.1 to the pairs of marginals $(\rho,\nu)$ and $(\rho,\nu^{n+1,n})$. We obtain
[…]
Invoking the known identities [47, Proposition 6.10]
$\mathrm{KL}(\nu^{n+1,n}|\nu)\leq\mathrm{KL}(\rho|\rho^{n,n})\,,\quad\mathrm{KL}(\nu|\nu^{n+1,n})\leq\mathrm{KL}(\nu|\nu^{n,n-1})\,,$
we arrive at the following bound
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})\leq\max\{1,\,\nicefrac{{\tau\Lambda}}{{\varepsilon}}\}(\mathrm{KL}(\rho|\rho^{n,n})+\mathrm{KL}(\nu|\nu^{n,n-1}))\,.$
Using this result in (2.7) gives
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})-\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n,n-1})\leq-\min\{1,\,\nicefrac{{\varepsilon}}{{\tau\Lambda}}\}\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})\,.$
We thus obtain from a simple recursion that for all $n\geq N-1$
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})\leq(1+\min\{1,\,\nicefrac{{\varepsilon}}{{\tau\Lambda}}\})^{-(n-N+2)}\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{N-1,N-2})\,.$
From the monotonicity bounds ([47, Proposition 6.5])
(2.8) $\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{m+1,m+1})\leq\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{m+1,m})\leq\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{m,m})\,,\quad\forall m\geq 0\,,$
we obtain
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})\leq(1+\min\{1,\,\nicefrac{{\varepsilon}}{{\tau\Lambda}}\})^{-(n-N+2)}\,\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{1,0})\,,\quad\forall n\geq N-1\,.$
The desired conclusion follows from the bounds
(2.9) $\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{1,0})\leq\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{0,0}),\quad\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n+1})\leq\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})\,,$
which are a consequence of (2.8).
Proof of (ii) Let now $n\geq N$. The standing assumptions grant that we can apply Theorem 1.1 to the pairs of marginals $(\rho,\nu)$ and $(\rho,\nu^{n,n-1})$. We thus obtain
(2.10) $\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n,n-1})\leq\mathrm{KL}(\nu|\nu^{n,n-1})+\frac{\Lambda}{2\varepsilon}\mathbf{W}^{2}_{2}(\nu^{n-1,n},\nu)\,.$
Since $\nu^{n,n-1}$ satisfies TI($\tau$) for $n\geq N$ we obtain
$\mathrm{KL}(\nu|\nu^{n,n-1})\geq\big{(}1+\nicefrac{{\tau\Lambda}}{{\varepsilon}}\big{)}^{-1}\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n,n-1})\,.$
Using this bound in (2.7) gives
$\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n+1,n})\leq\big{(}1+\nicefrac{{\varepsilon}}{{\tau\Lambda}}\big{)}^{-1}\,\mathrm{KL}(\pi_{\varepsilon}^{\rho\nu}|\pi^{n,n-1})\,.$
The desired conclusion follows from a simple recursion and the monotonicity bounds (2.8). ∎
Before moving to the computation of convergence rates, we recall that a probability measure $\rho$ satisfies a Poincaré inequality PI($C_{\rho}$) with constant $C_{\rho}>0$ if for any $f\in W^{1,2}(\rho)$ it holds
(PI($C_{\rho}$)) $\mathrm{Var}_{\rho}(f)\coloneqq\mathbb{E}_{\rho}[f^{2}(X)]-\mathbb{E}_{\rho}[f(X)]^{2}\leq C_{\rho}\,\int_{\mathcal{X}}|\nabla f|^{2}\mathrm{d}\rho\,.$
Moreover, let us recall here that any $\alpha$-log-concave measure satisfies the Talagrand transport inequality $\mathrm{TI}(\alpha^{-1})$ and the Poincaré inequality $\mathrm{P}(\alpha^{-1})$ (see [5, Corollaries 4.8.2 and 9.3.2]).
2.3. Proof of Proposition 1.4
Clearly we have $\nabla_{2}^{2}c(x,y)=\Sigma$ and hence
[…]
Therefore in this section, in order to study the semiconcavity of $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$, it is enough to control the conditional covariance matrices.
2.3.1. Log-concavity of $\rho$
We start with the proof of (ii). From (2.3) (applied along Sinkhorn’s algorithm) it follows $\nabla^{2}\varphi^{n}_{\varepsilon}(x)\succeq-\Sigma$ for any $n\in\mathbb{N}$, which combined with (1.6) further implies that for all $y$
$\nabla^{2}_{1}\big{(}-\log\pi^{n,n}(x|y)\big{)}\succeq\nabla^{2}U_{\rho}(y)\succeq\alpha_{\rho}\,,$
where we wrote $\pi^{n,n}(x|y)$ for the density of $\pi^{n,n}(\mathrm{d}x|y)$ with respect to the Lebesgue measure. This guarantees that, uniformly in $y\in\mathbb{R}^{d}$ and $n\in\mathbb{N}$, the conditional measure $\pi^{n,n}(\mathrm{d}x|y)$ satisfies the Poincaré inequality $\mathrm{PI}(\alpha_{\rho}^{-1})$ (cf. [5, Corollaries 4.8.2]).
Then, from (2.11) we deduce that for any unit vector $v$ it holds
$\varepsilon\,\langle v,\,\nabla_{2}^{2}(c(x,y)+\psi^{n}_{\varepsilon}(y))v\rangle=\mathrm{Var}_{X\sim\pi^{n,n}(\cdot|y)}(\langle v,\,\Sigma X\rangle)\leq\alpha_{\rho}^{-1}\,\|\Sigma v\|^{2}\leq\alpha_{\rho}^{-1}\,\|\Sigma\|_{2}^{2}\,.$
This means that for any $n\in\mathbb{N}$ and uniformly in $x\in\operatorname{supp}(\rho)$ the map $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave on $\operatorname{supp}(\nu)$ with $\Lambda=(\varepsilon\alpha_{\rho})^{-1}\|\Sigma\|_{2}^{2}$. This combined with Theorem 1.2-(i) proves the exponential proves convergence of Sinkhorn’s algorithm. If $\varepsilon\leq\|\Sigma\|_{2}\,\sqrt{\nicefrac{{\tau}}{{\alpha_{\rho}}}}$, then $\varepsilon\leq\tau\Lambda$ and the rate takes the form (1.11).
2.3.2. Log-concavity of $\rho$ and log-semiconvexity of $\nu$
We now discuss (i). Strengthening our assumption on the marginals, we can improve on the convergence rate. As a starter, we show the following convexity/concavity result that generalizes what is already known in for the Euclidean quadratic cost [20] and [12, Theorem 10].
[…]
Assume $\Sigma\succ 0$, $\nabla^{2}U_{\rho}\succeq\alpha_{\rho}$ and $\nabla^{2}U_{\nu}\preceq\beta_{\nu}$ for some $\alpha_{\rho}>0$ and $\beta_{\nu}\in(0,+\infty)$. If $\nabla^{2}\varphi^{0}_{\varepsilon}\succeq-\Sigma$, [^4] Usually Sinkhorn’s algorithm is initialised at $\varphi^{0}_{\varepsilon}=0$, for which this assumption holds. Nonetheless, as for any $n\geq 1$, $\nabla^{2}\varphi^{n}_{\varepsilon}\succeq-\Sigma$ by (2.3), one can get a similar conclusion for any initialization and all $n\geq 1$ by replacing (2.12) with $\nabla^{2}\varphi^{n}_{\varepsilon}\succeq-\Sigma+A_{n-1}\,\Sigma$ and $\nabla^{2}\psi^{n}_{\varepsilon}\preceq-\Sigma+B_{n-1}\,\Sigma$, where $(A_{n})_{n\in\mathbb{N}}$ and $(B_{n})_{n\in\mathbb{N}}$ satisfy (2.13). then for any $n\in\mathbb{N}$ we have
(2.12) $\nabla^{2}\varphi^{n}_{\varepsilon}\succeq-\Sigma+A_{n}\,\Sigma\quad\text{ and}\quad\nabla^{2}\psi^{n}_{\varepsilon}\preceq-\Sigma+B_{n}\,\Sigma\,,$
where $(A_{n})_{n\in\mathbb{N}}$ and $(B_{n})_{n\in\mathbb{N}}$ are two sequences of positive semidefinite symmetric matrices iteratively defined via
(2.13) $\begin{cases}B_{n}=\Sigma(A_{n}\Sigma+\varepsilon\alpha_{\rho})^{-1}\,,\\ A_{n+1}=\Sigma(B_{n}\Sigma+\varepsilon\beta_{\nu})^{-1}\,,\end{cases}\text{ started at }A_{0}=0\,,$
and converging to
(2.14) $A_{\infty}=$ $\,-\frac{\varepsilon\alpha_{\rho}}{2}\Sigma^{-1}+\biggl{(}\frac{\varepsilon^{2}\alpha_{\rho}^{2}}{4}\Sigma^{-2}+\frac{\alpha_{\rho}}{\beta_{\nu}}\biggr{)}^{\nicefrac{{1}}{{2}}}\,,$
$B_{\infty}=$ $\,-\frac{\varepsilon\beta_{\nu}}{2}\Sigma^{-1}+\biggl{(}\frac{\varepsilon^{2}\beta_{\nu}^{2}}{4}\Sigma^{-2}+\frac{\beta_{\nu}}{\alpha_{\rho}}\biggr{)}^{\nicefrac{{1}}{{2}}}\,.$
Proof.
The idea of the proof is to mimic the iterative proof given in [19, 12] combined with covariance bounds as in [20]. We will proceed by induction. The base case holds true with $A_{0}=0$, by assumption. Next, let us assume that $\nabla^{2}\varphi^{n}_{\varepsilon}\succeq-\Sigma+A_{n}\Sigma$ for some $A_{n}\succeq 0$. This implies that the conditional measure $\pi^{n,n}(\cdot|y)$ is log-concave since
$\nabla^{2}_{1}(-\log\pi^{n,n}(x|y))=\varepsilon^{-1}(\Sigma+\nabla^{2}\varphi^{n}_{\varepsilon}(x))+U_{\rho}\succeq\varepsilon^{-1}A_{n}\Sigma+\alpha_{\rho}\,.$
Then, from the Brascamp–-Lieb inequality (see for instance [20, Lemma 2]), we can deduce that
[…]
$\preceq-\Sigma+\Sigma(A_{n}\Sigma+\varepsilon\alpha_{\rho})^{-1}\Sigma\,,$
which proves $\nabla^{2}\psi^{n}_{\varepsilon}\preceq-\Sigma+B_{n}\Sigma$. This further implies that the conditional probability measure $\pi^{n+1,n}(\cdot|x)$ is log-convex uniformly in $x\in\mathbb{R}^{d}$ since
$\nabla^{2}_{2}(-\log\pi^{n+1,n}(y|x))=\varepsilon^{-1}(\Sigma+\nabla^{2}\psi^{n}_{\varepsilon}(y))+U_{\nu}(y)\preceq\varepsilon^{-1}\,B_{n}\Sigma+\beta_{\nu}\,.$
This, combined with the Cramér–Rao inequality (see for instance [20, Lemma 2]) gives uniformly in $x\in\mathbb{R}^{d}$
[…]
$\succeq-\Sigma+(B_{n}\Sigma+\varepsilon\beta_{\nu})^{-1}\Sigma\,,$
which proves $\nabla^{2}\varphi^{n+1}_{\varepsilon}\succeq-\Sigma+A_{n+1}\Sigma$. This shows the validity of (2.12) with the sequences defined at (2.13). Moreover, the same induction argument shows that whenever $\Sigma\succ 0$ we are guaranteed for any $n\in\mathbb{N}$ that $A_{n}\succ 0$, $B_{n}\succ 0$ and both matrices commute with $\Sigma$. The rest of the proof is devoted to showing that the sequence $(A_{n},B_{n})_{n\in\mathbb{N}}$ converge to the fixed point of (2.13), so that (2.14) follows from the validity of (2.13) in such limit points. In view of that, for any symmetric matrix $M$ let us denote $\Lambda(M)$ and $\lambda(M)$ respectively its biggest and smallest eigenvalue. Since $\Sigma$ is non-singular, notice that (2.13) can be rewritten as
$B_{n}=(A_{n}+\varepsilon\alpha_{\rho}\Sigma^{-1})^{-1}\,\quad\text{ and }\quad A_{n+1}=(B_{n}+\varepsilon\beta_{\nu}\Sigma^{-1})^{-1}\,.$
We will show that $(\lambda(A_{n}))_{n\in\mathbb{N}}$ is monotone increasing and $(\Lambda(B_{n}))_{n\in\mathbb{N}}$ is monotone decreasing. Clearly $\lambda(A_{1})-\lambda(A_{0})=\lambda(A_{1})>0$. Next, if we assume that $\lambda(A_{n})-\lambda(A_{n-1})\geq 0$ then we have
$\Lambda(B_{n})-\Lambda(B_{n-1})=(\lambda(A_{n}+\varepsilon\alpha_{\rho}\Sigma^{-1}))^{-1}-(\lambda(A_{n-1}+\varepsilon\alpha_{\rho}\Sigma^{-1}))^{-1}$
$=(\lambda(A_{n})+\varepsilon\alpha_{\rho}\Lambda(\Sigma)^{-1})^{-1}-(\lambda(A_{n-1})+\varepsilon\alpha_{\rho}\Lambda(\Sigma)^{-1})^{-1}<0\,,$
where we have used the fact that $A_{n}$ commutes with $\Sigma$, which implies that they are jointly diagonalizable and thus $\lambda(A_{n}+\varepsilon\alpha_{\rho}\Sigma^{-1})=\lambda(A_{n})+\varepsilon\alpha_{\rho}\lambda(\Sigma^{-1})$. From $\Lambda(B_{n})-\Lambda(B_{n-1})<0$, by reasoning analogously, we get
$\lambda(A_{n+1})-\lambda(A_{n})=(\Lambda(B_{n})+\varepsilon\beta_{\nu}\lambda(\Sigma)^{-1})^{-1}-(\Lambda(B_{n-1})+\varepsilon\beta_{\nu}\lambda(\Sigma)^{-1})>0\,.$
This shows that the sequence of positive semidefinite matrices $(B_{n})_{n\in\mathbb{N}}$ is compact in spectral norm, with norm strictly decreasing. Thus, the whole sequence $(B_{n})_{n\in\mathbb{N}}$ converges to a positive semidefinite matrix $B_{\infty}$ which solves
(2.15) $B_{\infty}=[(B_{\infty}+\varepsilon\beta_{\nu}\Sigma^{-1})^{-1}+\varepsilon\alpha_{\rho}\Sigma^{-1}]^{-1}\,,$
Observe that, since $B_{\infty}$ is positive semidefinite, the right hand side in the above identity dominates a positive definite matrix, whence $B_{\infty}$ is positive definite and invertible. Moreover, since $B_{\infty}$ and $\Sigma$ commute, we can rewrite (2.15) as
$0=$ $\,\alpha_{\rho}\,B_{\infty}^{2}+\varepsilon\alpha_{\rho}\beta_{\nu}B_{\infty}\Sigma^{-1}-\beta_{\nu}$
$=$ $\,\alpha_{\rho}\,\biggl{(}B_{\infty}+\frac{\varepsilon\beta_{\nu}}{2}\Sigma^{-1}+\biggl{(}\frac{\varepsilon^{2}\beta_{\nu}^{2}}{4}\Sigma^{-2}+\frac{\beta_{\nu}}{\alpha_{\rho}}\biggr{)}^{\nicefrac{{1}}{{2}}}\biggr{)}\biggl{(}B_{\infty}+\frac{\varepsilon\beta_{\nu}}{2}\Sigma^{-1}-\biggl{(}\frac{\varepsilon^{2}\beta_{\nu}^{2}}{4}\Sigma^{-2}+\frac{\beta_{\nu}}{\alpha_{\rho}}\biggr{)}^{\nicefrac{{1}}{{2}}}\biggr{)}\,,$
where the square root has to be understood as the square root of a positive semidefinite matrix. Since $B_{\infty}$ ought to be positive definite we conclude that
$B_{\infty}=-\frac{\varepsilon\beta_{\nu}}{2}\Sigma^{-1}+\biggl{(}\frac{\varepsilon^{2}\beta_{\nu}^{2}}{4}\Sigma^{-2}+\frac{\beta_{\nu}}{\alpha_{\rho}}\biggr{)}^{\nicefrac{{1}}{{2}}}\,.$
The convergence of $(B_{n})_{n\in\mathbb{N}}$ to $B_{\infty}$ implies the convergence of $(A_{n})_{n\in\mathbb{N}}$ towards the limit matrix
$A_{\infty}=(B_{\infty}+\varepsilon\beta_{\nu}\Sigma^{-1})^{-1}=-\frac{\varepsilon\alpha_{\rho}}{2}\Sigma^{-1}+\biggl{(}\frac{\varepsilon^{2}\alpha_{\rho}^{2}}{4}\Sigma^{-2}+\frac{\alpha_{\rho}}{\beta_{\nu}}\biggr{)}^{\nicefrac{{1}}{{2}}}\,.$
If in the previous lemma we consider as Sinkhorn’s iterates the constant iterates $\varphi^{n}_{\varepsilon}=\varphi^{\nu}_{\varepsilon}$ and $\psi^{n}_{\varepsilon}=\psi^{\nu}_{\varepsilon}$ (i.e. we start the algorithm already in the fixed point $(\varphi^{\nu}_{\varepsilon},\psi^{\nu}_{\varepsilon})$), then we immediately deduce the following.
[…]
When the two marginals satisfy (1.9) and $\Sigma\succ 0$, we have
$\nabla^{2}\varphi^{\nu}_{\varepsilon}$ $\succeq-\Sigma-\frac{\varepsilon\alpha_{\rho}}{2}+\biggl{(}\frac{\varepsilon^{2}\alpha_{\rho}^{2}}{4}+\frac{\alpha_{\rho}}{\beta_{\nu}}\,\Sigma^{2}\biggr{)}^{\nicefrac{{1}}{{2}}}\,,$
$\nabla^{2}\psi^{\nu}_{\varepsilon}$ $\preceq-\Sigma-\frac{\varepsilon\beta_{\nu}}{2}+\biggl{(}\frac{\varepsilon^{2}\beta_{\nu}^{2}}{4}+\frac{\beta_{\nu}}{\alpha_{\rho}}\,\Sigma^{2}\biggr{)}^{\nicefrac{{1}}{{2}}}\,.$
Let us also highlight here a regularizing property that can be immediately deduced from Section 2.3.2.
[…]
When the two marginals satisfy (1.9) and $\Sigma\succ 0$, if $A_{0}\succeq A_{\infty}$ (defined at (2.14)) then $B_{n}\preceq B_{\infty}$ for any $n\in\mathbb{N}$. Equivalently, if we start Sinkhorn’s algorithm in $\varphi^{0}_{\varepsilon}$ such that
(2.16) $\nabla^{2}\varphi^{0}_{\varepsilon}\succeq-\Sigma-\frac{\varepsilon\alpha_{\rho}}{2}+\biggl{(}\frac{\varepsilon^{2}\alpha_{\rho}^{2}}{4}+\frac{\alpha_{\rho}}{\beta_{\nu}}\,\Sigma^{2}\biggr{)}^{\nicefrac{{1}}{{2}}}\,,$
then for any $n\in\mathbb{N}$
(2.17) $\nabla^{2}\psi^{n}_{\varepsilon}\preceq-\Sigma-\frac{\varepsilon\beta_{\nu}}{2}+\biggl{(}\frac{\varepsilon^{2}\beta_{\nu}^{2}}{4}+\frac{\beta_{\nu}}{\alpha_{\rho}}\,\Sigma^{2}\biggr{)}^{\nicefrac{{1}}{{2}}}\,.$
Proof.
Let us start by showing that if $A_{0}\succeq A_{\infty}$, then $B_{n}\preceq B_{\infty}$. In view of this, it suffices to show by induction that $A_{n}-A_{\infty}\succeq 0$ and $B_{n}-B_{\infty}\preceq 0$. The base case is the standing assumption. Next, assume $A_{n}-A_{\infty}\succeq 0$. From (2.13) (and by recalling that $(A_{\infty},B_{\infty})$ is a fixed point solution of the latter) we immediately deduce that
$B_{n}-B_{\infty}=(A_{n}+\varepsilon\alpha_{\rho}\Sigma^{-1})^{-1}-(A_{\infty}+\varepsilon\alpha_{\rho}\Sigma^{-1})^{-1}\preceq 0\,,$
since our inductive step implies $(A_{n}+\varepsilon\alpha_{\rho}\Sigma^{-1})-(A_{\infty}+\varepsilon\alpha_{\rho}\Sigma^{-1})\succeq 0$. This shows $B_{n}-B_{\infty}\preceq 0$, which further implies (same reasoning from (2.13)) $A_{n+1}-A_{\infty}\succeq 0$, which concludes our inductive proof.
Finally, if (2.16) holds, then we can take $A_{0}\succeq A_{\infty}$ in Section 2.3.2, hence the above discussion yields $B_{n}\preceq B_{\infty}$ which implies (2.17). ∎
We are now able to conclude our analysis of the semiconcavity of $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$. For a generic initial condition, we know by Lemma 2.3.2 that $\nabla\psi^{n}_{\varepsilon}\preceq-\Sigma+B_{n}\Sigma$. Therefore for any $n\in\mathbb{N}$ the map $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\left\|B_{n}\,\Sigma\right\|_{2}$-semiconcave. Since $B_{n}$ converges to $B_{\infty}$ we have
$\lim_{n\to\infty}\left\|B_{n}\,\Sigma\right\|_{2}=\left\|B_{\infty}\Sigma\right\|_{2}=-\frac{\varepsilon\beta_{\nu}}{2}+\biggl{(}\frac{\varepsilon^{2}\beta_{\nu}^{2}}{4}+\frac{\beta_{\nu}}{\alpha_{\rho}}\Sigma^{2}\biggr{)}^{\nicefrac{{1}}{{2}}}$
$\leq-\frac{\varepsilon\beta_{\nu}}{2}+\biggl{(}\frac{\varepsilon^{2}\beta_{\nu}^{2}}{4}+\frac{\beta_{\nu}}{\alpha_{\rho}}\left\|\Sigma\right\|_{2}^{2}\biggr{)}^{\nicefrac{{1}}{{2}}}\leq\sqrt{\nicefrac{{\beta_{\nu}}}{{\alpha_{\rho}}}}\,\left\|\Sigma\right\|_{2}\,,$
and therefore Theorem 1.2-(i) implies the exponential convergence of Sinkhorn’s algorithm and for $\varepsilon\leq\tau\,\sqrt{\nicefrac{{\beta_{\nu}}}{{\alpha_{\rho}}}}\,\left\|\Sigma\right\|_{2}$ the asymptotic rate is
(2.18) $1-\frac{\varepsilon}{\varepsilon+\tau\|\Sigma\|_{2}(\nicefrac{{\beta_{\nu}}}{{\alpha_{\rho}}})^{\nicefrac{{1}}{{2}}}}\,.$
Finally, if we start Sinkhorn’s algorithm with $\varphi^{0}_{\varepsilon}$ satisfying (2.16), then Section 2.3.2 implies the validity of (2.17) for all $n\in\mathbb{N}$. Thus, $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave for all $n\in\mathbb{N}$ with $\Lambda=\left\|B_{\infty}\Sigma\right\|_{2}\leq\sqrt{\nicefrac{{\beta_{\nu}}}{{\alpha_{\rho}}}}\|\Sigma\|_{2}$. Therefore in this case Theorem 1.2-(i) gives the exponential convergence of Sinkhorn’s algorithm with rate given by (2.18) if $\varepsilon\leq\tau\,\sqrt{\nicefrac{{\beta_{\nu}}}{{\alpha_{\rho}}}}\,\left\|\Sigma\right\|_{2}$.
2.3.3. Marginal $\rho$ weakly log-concave
Here we restrict to the quadratic cost setting (i.e., $\Sigma=\operatorname{Id}$) and relax the convexity assumption of $U_{\rho}$ by simply requiring $\rho$ to be weakly log-concave (cf. (1.12)), that is, we assume
(2.19) $\langle\nabla U_{\rho}(\hat{x})-\nabla U_{\rho}(x),\hat{x}-x\rangle\geq\alpha_{\rho}|\hat{x}-x|^{2}-|\hat{x}-x|f_{L}(|\hat{x}-x|)\,,\quad\forall x,\hat{x}\in\mathbb{R}^{d},$
where for any $L\geq 0$, $r>0$, $f_{L}(r)=2L^{1/2}\tanh\big{(}(L^{1/2}r)/2\big{)}$. We now reformulate this condition introducing the convexity profile (this is a classical notion in the coupling literature, see [27] for example) of a given $U$ as the function $\kappa_{U}:\mathbb{R}_{+}\longrightarrow\mathbb{R}$ given by
(2.20) $\kappa_{U}(r)\coloneqq\inf\biggl{\{}\frac{\langle\nabla U(\hat{x})-\nabla U(y),\,x-\hat{x}\rangle}{|x-\hat{x}|^{2}}\quad\colon\,|x-\hat{x}|=r\biggr{\}}\,.$
For ease of notation we also set $\kappa_{\rho}\coloneqq\kappa_{U_{\rho}}$. With this notation at hand, (2.19) rewrites as
(2.21) $\kappa_{\rho}(r)\geq\alpha_{\rho}-r^{-1}f_{L}(r)\quad\forall r>0\,.$
It follows from (2.3) that, since $\nabla^{2}_{1}c=\operatorname{Id}$,
(2.22) $\kappa_{(\varepsilon^{-1}\varphi^{n}_{\varepsilon}+U_{\rho})}(r)\geq\alpha_{\rho}-\varepsilon^{-1}-r^{-1}f_{L}(r)\,\quad\forall r>0,n\in\mathbb{N}\,.$
At this point, we can invoke [19, Lemma 3.1] (see Appendix A for a statement of this result using the current notation) to conclude that
$\kappa_{\Phi^{\rho}_{0}(\varphi^{n}_{\varepsilon})}(r)\geq\frac{\varepsilon\alpha_{\rho}-1}{\varepsilon\alpha_{\rho}}-\frac{L}{\varepsilon\alpha_{\rho}^{2}}\,,$
which gives
$\nabla^{2}\psi^{n}_{\varepsilon}=-\nabla^{2}\Phi^{\rho}_{0}(\varphi^{n}_{\varepsilon})\preceq-\frac{\varepsilon\alpha_{\rho}-1}{\varepsilon\alpha_{\rho}}+\frac{L}{\varepsilon\alpha_{\rho}^{2}}=-1+\frac{1}{\varepsilon\alpha_{\rho}}+\frac{L}{\varepsilon\alpha_{\rho}^{2}}\,.$
From this we immediately conclude that uniformly in $x\in\operatorname{supp}(\rho)$, for any $n\in\mathbb{N}$, the map $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave with $\Lambda=(\varepsilon\alpha_{\rho})^{-1}+\nicefrac{{L}}{{\varepsilon\alpha_{\rho}^{2}}}$. We can thus invoke Theorem 1.2-(i) which gives exponential convergence of Sinkhorn’s algorithm. Moreover, if $\varepsilon\leq\alpha_{\rho}^{-1}\sqrt{\tau(\alpha_{\rho}+L)}$, then the convergence rate is
$1-\frac{\varepsilon^{2}\alpha_{\rho}^{2}}{\varepsilon^{2}\alpha_{\rho}^{2}+\tau(\alpha_{\rho}+L)}\,.$
$\square$
2.4. Proof of Proposition 1.4
Assume that the cost satisfies
$\nabla_{1}^{2}c(x,y)\,,\nabla^{2}_{2}c(x,y)\succeq h(c)\quad\text{ and }\quad\left\|\nabla_{12}^{2}c(x,y)\right\|_{2}\vee\left\|\nabla_{1}^{2}c(x,y)\right\|_{2}\leq H(c)\,,$
uniformly in $x,y\in\mathbb{R}^{d}$. Moreover assume that $\rho$ has light tails in the following sense
$\nabla^{2}U_{\rho}(x)\succeq\begin{cases}C\,|x|^{\delta}\quad&\text{ for }|x|>R\,,\\ -L\quad&\text{ for }|x|\leq R\,,\end{cases}$
for some positive $\delta,\,C>0$ and $L,\,R\geq 0$. To estimate the conditional covariances, we shall use the following abstract result.
[…]
Let $\pi=e^{-U}$, $\alpha>0,L,R\geq 0$ be such that
(2.23) $\nabla^{2}U(x)\succeq\begin{cases}-L,\quad&{|x|<R}\,,\\ \alpha,\quad&|x|\geq R\,.\end{cases}$
Moreover, let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ be Lipschitz . Then we have
$\mathrm{Cov}_{X\sim\pi}(f(X))\preceq 2\|f\|^{2}_{\mathrm{Lip}}\big{(}\alpha^{-1}+\alpha^{-2}(L+\alpha)^{2}R^{2}\big{)}\,.$
Proof.
We define the function $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$ given by
$\phi(x)=\frac{L+\alpha}{2}|x|^{2}\mathbf{1}_{\{|x|<R\}}+\Big{(}(L+\alpha)R(|x|-R)+\frac{(L+\alpha)R^{2}}{2}\Big{)}\mathbf{1}_{\{|x|\geq R\}}\,.$
Then, the function $\phi$ is convex and $(L+\alpha)R$-Lipschitz. Moreover, $\nabla^{2}\phi(x)\succeq(L+\alpha)$ for $|x|<R$. But then, we deduce from (2.23) that $\bar{\pi}(\mathrm{d}x)\propto\exp(-U(x)-\phi(x))\mathrm{d}x$ is $\alpha$-log-concave and in particular satisfies $\mathrm{PI}(\alpha^{-1})$. Now, let $(X,\bar{X})$ be a coupling between $\pi$ and $\bar{\pi}$. Then, following the argument given in [8, Lemma 2.1], we find that for all $v$ s.t. $|v|=1$
(2.24) $\langle v,\,\mathrm{Cov}_{X\sim\pi}(f(X))v\rangle$ $=\mathrm{Var}_{X\sim\pi}(\langle f(X),v\rangle)$
$\leq 2\mathrm{Var}_{\bar{X}\sim\bar{\pi}}(\langle f(\bar{X}),v\rangle)+2\mathbb{E}[|\langle f(X),v\rangle-\langle f(\bar{X}),v\rangle|^{2}]\,.$
Using that the function $x\mapsto\langle f(x),v\rangle$ is $\|f\|_{\mathrm{Lip}}$-Lipschitz, we can use that $\bar{\pi}$ satisfies $\mathrm{PI}(\alpha^{-1})$ to obtain
$\mathrm{Var}_{\bar{X}\sim\bar{\pi}}(\langle f(\bar{X}),v\rangle)\leq\|f\|^{2}_{\mathrm{Lip}}\alpha^{-1}\,.$
Moreover, if $(X,\bar{X})$ is optimal for $\mathbf{W}_{2}(\pi,\bar{\pi})$ we obtain
$\mathbb{E}[|\langle f(X),v\rangle-\langle f(\bar{X}),v\rangle|^{2}]\leq\|f\|^{2}_{\mathrm{Lip}}\mathbb{E}[{\sf d}(X,\bar{X})^{2}]=\|f\|^{2}_{\mathrm{Lip}}\mathbf{W}_{2}^{2}(\pi,\hat{\pi})\,.$
Plugging the last two bounds back into (2.24) and invoking first TI($\alpha^{-1}$) then LSI($\alpha^{-1}$) for $\bar{\pi}$ we obtain
$\langle v,\,\mathrm{Cov}_{f(X)\sim\pi}(f(X)),v\rangle\leq 2\|f\|^{2}_{\mathrm{Lip}}\big{(}\alpha^{-1}+\alpha^{-2}\|\phi\|^{2}_{\mathrm{Lip}}\big{)}\,.$
Since $\|\phi\|_{\mathrm{Lip}}\leq(L+\alpha)R$, the conclusion follows. ∎
With this result at hand, we obtain the following.
[…]
Assume that $c$ satisfies (1.15) and that $\rho$ satisfies (1.14). Then for any $\alpha>0$ we have
(2.25) $\sup_{y\in\mathcal{Y}}\left\|\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\right\|_{2}\leq 2H(c)^{2}\,\biggl{(}\frac{\varepsilon}{\alpha}+\biggl{(}\frac{L\varepsilon}{\alpha}+\frac{\alpha+\delta_{H}}{\alpha}\biggr{)}\,\biggl{[}R^{2}\vee\biggl{(}\frac{\alpha+\delta_{H}}{\varepsilon\,C}\biggr{)}^{\frac{2}{\delta}}\biggr{]}\biggr{)}\,,$
where we have set $\delta_{H}\coloneqq H(c)-h(c)$. As a consequence of this we have
$\sup_{y\in\mathcal{Y}}\left\|\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\right\|_{2}\leq 2H(c)^{2}\,\biggl{(}1+(L+2)\biggl{[}R^{2}\vee{C^{-\nicefrac{{2}}{{\delta}}}}\biggl{(}1+\frac{\delta_{H}}{\varepsilon}\biggr{)}^{\nicefrac{{2}}{{\delta}}}\biggr{]}\biggr{)}\,.$
Proof.
Notice that the upper bound on the Hessian of the cost guarantees the semiconvexity of $\varphi^{n}_{\varepsilon}$, since we recall from (2.3) that
$\nabla^{2}\varphi^{n}_{\varepsilon}(x)=-\int_{\mathcal{Y}}\nabla_{1}^{2}c(x,y)\,\pi^{n,n-1}(\mathrm{d}y|x)+\varepsilon^{-1}\,\mathrm{Cov}_{Y\sim\pi^{n,n-1}(\cdot|x)}(\nabla_{1}c(x,Y))\succeq-H(c)\,.$
By using this estimate we deduce that uniformly in $y\in\mathcal{Y}$ it holds
$\nabla^{2}_{1}(-\log\pi^{n,n}(x|y))\succeq\begin{cases}C\,|x|^{\delta}-\frac{\delta_{H}}{\varepsilon}\quad&\text{ for }|x|>R\,,\\ -L-\frac{\delta_{H}}{\varepsilon}\quad&\text{ for }|x|\leq R\,.\end{cases}$
Therefore, for any fixed $\alpha>0$, if we set
$R(\alpha,\varepsilon)=R\vee\biggl{(}\frac{\alpha+\delta_{H}}{\varepsilon\,C}\biggr{)}^{\frac{1}{\delta}}\,,$
we find
$\nabla^{2}_{1}(-\log\pi^{n,n}(x|y))\succeq\begin{cases}\frac{\alpha}{\varepsilon}\quad&\text{ for }|x|>R(\alpha,\varepsilon)\,,\\ -L-\frac{\delta_{H}}{\varepsilon}\quad&\text{ for }|x|\leq R(\alpha,\varepsilon)\,.\end{cases}$
We can now apply Section 2.4 to $\pi^{n,n}(\cdot|y)$ with $f(x)=\nabla_{2}c(x,y)$ (for which $\left\|f\right\|_{\mathrm{Lip}}\leq H(c)$) to obtain that, uniformly in $y\in\mathbb{R}^{d}$, it holds
$\mathrm{Cov}_{X\sim\pi}(\nabla_{2}c(X,y))\preceq 2H(c)^{2}\,\biggl{(}\frac{\varepsilon}{\alpha}+\biggl{(}\frac{L\varepsilon+\delta_{H}}{\alpha}+1\biggr{)}\,\biggl{[}R^{2}\vee\biggl{(}\frac{\alpha+\delta_{H}}{\varepsilon\,C}\biggr{)}^{\frac{2}{\delta}}\biggr{]}\biggr{)}\,.$
We have thus shown the first bound. The second one can be obtained by considering the specific choice $\bar{\alpha}=\varepsilon+\delta_{H}$. ∎
As a corollary of the previous estimate we finally deduce the convergence rate for Sinkhorn’s algorithm in the light-tails regime.
[…]
Assume that $c$ satisfies (1.15) and that $\rho$ satisfies (1.14). Then Sinkhorn’s algorithm converges exponentially fast. Moreover, if $H(c)=h(c)$ then, for
(2.26) $\varepsilon\leq\sqrt{2\tau H(c)^{2}\,(1+(L+2)\,[R^{2}\vee C^{-\nicefrac{{2}}{{\delta}}}])}$
the rate is
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\tau H(c)^{2}(1+(L+2)\,[R^{2}\vee C^{-\nicefrac{{2}}{{\delta}}}])}\,.$
Otherwise, if $H(c)>h(c)$ then, for
(2.27) $\varepsilon\leq 1\wedge\frac{\delta_{H}}{R^{\delta}C-1}\wedge\biggl{(}\tau\delta_{H}+2\tau H(c)^{2}\,(1+(L+2)\,C^{-\nicefrac{{2}}{{\delta}}}(1+\delta_{H})^{\nicefrac{{2}}{{\delta}}}\biggr{)}^{\nicefrac{{\delta}}{{2+2\delta}}}$
we have rate of convergence
$1-\frac{\varepsilon^{2+\nicefrac{{2}}{{\delta}}}}{\varepsilon^{2}+\tau\varepsilon\delta_{H}+2\tau H(c)^{2}\,(1+(L+2)\,C^{-\nicefrac{{2}}{{\delta}}}(\varepsilon+\delta_{H})^{\nicefrac{{2}}{{\delta}}})}\,.$
Proof.
From Section 2.4 and (2.2) we deduce that
$\nabla^{2}_{2}(c(x,y)+\psi^{\nu}_{\varepsilon}(y))\preceq$ $\,H(c)-h(c)+\varepsilon^{-1}\sup_{y\in\mathcal{Y}}\left\|\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\right\|_{2}$
$\leq$ $\,\delta_{H}+\frac{2H(c)^{2}}{\varepsilon}\,\biggl{(}1+(L+2)\biggl{[}R^{2}\vee{C^{-\nicefrac{{2}}{{\delta}}}}\biggl{(}1+\frac{\delta_{H}}{\varepsilon}\biggr{)}^{\nicefrac{{2}}{{\delta}}}\biggr{]}\biggr{)}\eqqcolon\Lambda\,.$
This proves that $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave uniformly in $n\in\mathbb{N}$ and in $x\in\mathbb{R}^{d}$. Thus, from Theorem 1.2 we deduce the convergence of Sinkhorn’s algorithm and its rate. On the one hand, when $\delta_{H}=H(c)-h(c)=0$, for $\varepsilon$ small enough as in (2.26), the rate is
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\tau H(c)^{2}\,(1+(L+2)\,[R^{2}\vee C^{-\nicefrac{{2}}{{\delta}}}])}\,.$
On the other hand, if $H(c)>h(c)$ then, eventually for $\varepsilon$ small enough, we have
$R^{2}\leq C^{-\nicefrac{{2}}{{\delta}}}\biggl{(}1+\frac{\delta_{H}}{\varepsilon}\biggr{)}^{\nicefrac{{2}}{{\delta}}}$
and hence Theorem 1.2 implies the convergence of Sinkhorn’s algorithm and for $\varepsilon$ satisfying (2.27) the rate is
$1-\frac{\varepsilon^{2+\nicefrac{{2}}{{\delta}}}}{\varepsilon^{2+\nicefrac{{2}}{{\delta}}}+\varepsilon^{1+\nicefrac{{2}}{{\delta}}}\tau\delta_{H}+2\tau H(c)^{2}\,(\varepsilon^{\nicefrac{{2}}{{\delta}}}+(L+2)\,C^{-\nicefrac{{2}}{{\delta}}}(\varepsilon+\delta_{H})^{\nicefrac{{2}}{{\delta}}})}\,.$
2.4.1. Quadratic cost
For the Euclidean quadratic cost, the previous discussion (with $H(c)=h(c)=1$) combined with Theorem 1.2 shows the exponential convergence of Sinkhorn’s algorithm when the tails of $\rho$ are light. More precisely, if (1.14) holds for some $C,\,\delta>0$ and $R,\,L\geq 0$, then Sinkhorn’s algorithm exponential converges with rate $1-\Theta(\nicefrac{{\varepsilon^{2}}}{{\tau(1+L\,[R^{2}\vee C^{-\nicefrac{{2}}{{\delta}}}])}})$ as $\varepsilon\rightarrow 0$, see Table 1 for the meaning of $\Theta(\cdot)$.
2.4.2. Anisotropic quadratic costs and subspace elastic costs
As we have already noticed in (2.11), when considering the anisotropic quadratic cost $c(x,y)\coloneqq\langle x-y,\,\Sigma(x-y)\rangle/2$ we have
[…]
and hence it is enough to bound uniformly in $y\in\mathbb{R}^{d}$, for any unit vector $v$, the variance $\mathrm{Var}_{X\sim\pi^{n,n}(\cdot|y)}(\langle v,\,\Sigma X\rangle)$. This can be done via Poincaré inequality, as explained above with $\rho$ satisfying the light-tails condition (1.14). More precisely, we can reason as in Section 2.4 to deduce that
$\varepsilon\,\langle v,\,\nabla_{2}^{2}(c(x,y)+\psi^{n}_{\varepsilon}(y))v\rangle=\,\mathrm{Var}_{\pi^{n,n}(\cdot|y)}(\langle v,\Sigma\,X\rangle)\leq 2\|\Sigma\|_{2}^{2}\,\bigl{(}1+(L+2)\bigl{[}R^{2}\vee{C^{-\nicefrac{{2}}{{\delta}}}}\bigr{]}\bigr{)}\,.$
From this, we conclude that $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave uniformly in $n\in\mathbb{N}$ and in $x\in\mathbb{R}^{d}$, with $\Lambda=2\|\Sigma\|_{2}^{2}\,(1+(L+2)[R^{2}\vee{C^{-\nicefrac{{2}}{{\delta}}}}])$, which combined with our main result Theorem 1.2 implies the exponential convergence of Sinkhorn’s algorithm and for
(2.28) $\varepsilon\leq\left\|\Sigma\right\|_{2}\,\sqrt{2\,\tau(1+(L+2)[R^{2}\vee{C^{-\nicefrac{{2}}{{\delta}}}}])}$
the rate is
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\,\tau\,\|\Sigma\|_{2}^{2}\,\bigl{(}1+(L+2)\bigl{[}R^{2}\vee{C^{-\nicefrac{{2}}{{\delta}}}}\bigr{]}\bigr{)}}\,.$
2.5. Proof of Proposition 1.4
Assume that
$h(c)\preceq\nabla^{2}_{2}c(x,y)\preceq H(c)\quad\text{ and }\quad\nabla^{2}_{1}c(x,y)\preceq H(c)\,.$
We distinguish two cases.
2.5.1. Cost Lipschitz w.r.t. $y\in\mathcal{Y}$
If the cost is Lipschitz in $y$, uniformly in $x$, that is,
$\sup_{x\in\mathbb{R}^{d}}|c(x,y)-c(x,\hat{y})|\leq\mathrm{Lip}_{\infty,2}(c)\,|y-\hat{y}|\quad\forall y,\hat{y}\in\mathbb{R}^{d}\,,$
then we deduce from (2.2) that
$\nabla^{2}_{2}(c(x,y)+\psi^{n}_{\varepsilon}(y))\preceq H(c)-h(c)+\varepsilon^{-1}\,\mathrm{Lip}_{\infty,2}^{2}(c)\,.$
Therefore $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave uniformly in $x\in\mathbb{R}^{d}$ and in $n\in\mathbb{N}$ with $\Lambda=H(c)-h(c)+\varepsilon^{-1}\,\mathrm{Lip}^{2}_{\infty,2}(c)$. Theorem 1.2 implies then the convergence of Sinkhorn’s algorithm and for
(2.29) $\varepsilon\leq\frac{\tau(H(c)-h(c))+\sqrt{\tau^{2}(H(c)-h(c))^{2}+4\tau\mathrm{Lip}_{\infty,2}(c)^{2}}}{2}$
the rate is
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+\varepsilon\tau(H(c)-h(c))+\tau\mathrm{Lip}_{\infty,2}(c)^{2}}\,.$
2.5.2. Cost Lipschitz w.r.t. $x\in\mathcal{X}$
Assume our cost is Lipschitz in $x$, uniformly in $y$, i.e., such that
$\sup_{y\in\mathbb{R}^{d}}|c(x,y)-c(\hat{x},y)|\leq\mathrm{Lip}_{1,\infty}(c)\,|x-\hat{x}|\quad\forall x,\hat{x}\in\mathbb{R}^{d}.$
Moreover, we further assume that $\left\|\nabla_{12}^{2}c(x,y)\right\|_{2}\leq H(c)$ and that $\rho$ satisfies LSI($C_{\rho}$).
Notice that the Lipschitz continuity of the cost directly propagates to the Sinkhorn iterate $\varphi^{n}_{\varepsilon}$ since $\nabla\varphi^{n}_{\varepsilon}(x)=-\int\nabla_{1}c(x,y)\,\pi^{n,n-1}(\mathrm{d}y|x)$ and hence $\mathrm{Lip}(\varphi^{n}_{\varepsilon})\leq\mathrm{Lip}_{1,\infty}$, see also [25, Prop 2.4]. This key observation allows to prove the following lemma.
[…]
Assume $c$ satisfies (1.17) and that $\left\|\nabla_{12}^{2}c(x,y)\right\|_{2}\vee\left\|\nabla_{1}^{2}c(x,y)\right\|_{2}\leq H(c)$ and further that $\rho$ satisfies $LSI(C_{\rho})$. Then
$\sup_{y\in\mathcal{Y}}\left\|\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\right\|_{2}\leq 2H(c)^{2}C_{\rho}\,(1+\nicefrac{{4C_{\rho}\mathrm{Lip}_{1,\infty}^{2}}}{{\varepsilon^{2}}})\,.$
Proof.
The proof resembles to that of Section 2.4. Arguing as we did there, for any $v$ with $|v|=1$ the Poincaré inequality PI($C_{\rho}$) and the Talagrand $\mathrm{TI}(C_{\rho})$ combined with the $\mathrm{LSI}(C_{\rho})$ yield
$\langle v,\,\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\,v\rangle\leq$  $\,2H(c)^{2}\,\mathbf{W}_{2}^{2}(\pi^{n,n}(\cdot|y),\rho)+2\,\mathrm{Var}_{\hat{X}\sim\rho}(\langle v,\,\nabla_{2}c(\hat{X},y)\rangle)$
$\leq$  $\,4H(c)^{2}\,C_{\rho}\,\mathrm{KL}(\pi^{n,n}(\cdot|y)|\rho)+2C_{\rho}\int|\nabla_{12}^{2}c(x,y)v|^{2}\mathrm{d}\rho$
$\leq$  $\,2H(c)^{2}C_{\rho}\biggl{(}\frac{C_{\rho}}{\varepsilon^{2}}\int|\nabla_{1}(c(x,y)+\varphi^{n}_{\varepsilon}(x))|^{2}\pi^{n,n}(\mathrm{d}x|y)+1\biggr{)}$
$\leq$  $\,2H(c)^{2}C_{\rho}\,(1+\nicefrac{{4C_{\rho}\mathrm{Lip}^{2}_{1,\infty}}}{{\varepsilon^{2}}})\,.$
As a corollary of the previous estimate, we finally deduce the convergence rate for Sinkhorn’s algorithm in this setting. In fact, it is enough to notice that from (2.2) we have
$\nabla^{2}_{2}(c(x,y)+\psi^{n}_{\varepsilon}(y))\preceq$ $\,H(c)-h(c)+\varepsilon^{-1}\,\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))$
$\preceq$ $\,H(c)-h(c)+\frac{2H(c)^{2}\,C_{\rho}}{\varepsilon}\,(1+\nicefrac{{4C_{\rho}\mathrm{Lip}_{1,\infty}^{2}}}{{\varepsilon^{2}}})\,.$
Therefore, the map $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave uniformly in $x\in\mathbb{R}^{d}$ and in $n\in\mathbb{N}$ with $\Lambda=H(c)-h(c)+\frac{2H(c)^{2}\,C_{\rho}}{\varepsilon}\,(1+\nicefrac{{4C_{\rho}\mathrm{Lip}_{1,\infty}^{2}}}{{\varepsilon^{2}}})$. Theorem 1.2 implies then the convergence of Sinkhorn’s algorithm and for
(2.30) $\varepsilon\leq 1\wedge\biggl{(}\tau(H(c)-h(c))+2\tau H(c)^{2}C_{\rho}(1+4C_{\rho}\mathrm{Lip}_{1,\infty}^{2})\biggr{)}^{\nicefrac{{1}}{{4}}}$
the rate is
$1-\frac{\varepsilon^{4}}{\varepsilon^{4}+\tau\varepsilon^{3}(H(c)-h(c))+2\tau H(c)^{2}C_{\rho}(\varepsilon^{2}+4C_{\rho}\mathrm{Lip}_{1,\infty}^{2})}\,.$
$\square$
2.6. Proof of Proposition 1.4
Whenever we assume that $\rho$ is compactly supported, we aim to estimate the $\Lambda$-semiconcavity of $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ by relying on (2.2) and the fact that the covariance under the conditional probability $\pi^{n,n}(\cdot|y)$ has bounded support, i.e. $\operatorname{supp}(\pi^{n,n}(\cdot|y))\subseteq\operatorname{supp}(\rho)$. In order to do that we need an additional assumption either on the cost function or on the second marginal $\nu$.
2.6.1. Case $\nabla_{2}^{2}c(x,y)=\Sigma$
Under this additional assumption, (2.2) reads as
$\nabla^{2}(c(x,y)+\psi^{n}_{\varepsilon}(y))=\varepsilon^{-1}\,\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\,,$
which means that we can compute the semiconcavity parameter $\Lambda$ by estimating
$\sup_{y\in\mathcal{Y}}\left\|\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\right\|_{2}\,.$
To bound this, note that $\nabla_{2}^{2}c(x,y)=\Sigma$ implies that $\nabla_{2}c(x,y)=\Sigma\,y+g(x)$ is affine w.r.t. $y$ and hence
$\sup_{y\in\mathcal{Y}}\left\|\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(g(X))\right\|_{2}\leq\max_{x\in\operatorname{supp}(\rho)}\left\|g(x)\right\|_{2}^{2}=\left\|g\right\|_{\mathrm{L}^{\infty}(\rho)}^{2}\,.$
This implies that, uniformly in $x\in\operatorname{supp}(\rho)$ for every $n\in\mathbb{N}$, the function $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave with parameter $\Lambda\leq\varepsilon^{-1}\left\|g\right\|_{\mathrm{L}^{\infty}(\rho)}^{2}$, which combined with Theorem 1.2 implies the exponential convergence of Sinkhorn’s algorithm and for $\varepsilon\leq\sqrt{\tau}\,\left\|g\right\|_{\mathrm{L}^{\infty}(\rho)}$ the rate is
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+\tau\|g\|_{\mathrm{L}^{\infty}(\rho)}^{2}}\,.$
In particular, note that the convergence rate proven here is independent of the matrix $\Sigma$.
This setting covers the quadratic cost $c(x,y)=|x-y|^{2}/2$ over $\mathbb{R}^{d}$, where $\Sigma=\operatorname{Id}$, $\nabla_{2}c(x,y)=y-x$ and hence for any $\rho$ such that $\operatorname{supp}\rho\subseteq B_{R}(0)$ it holds $\Lambda\leq\nicefrac{{R^{2}}}{{\varepsilon}}$. Thus, the convergence rate of Sinkhorn’s algorithm, for $\varepsilon$ small enough, is $1-\Theta(\nicefrac{{\varepsilon^{2}}}{{\tau R^{2}}})$, see Table 1 for the meaning of $\Theta$. Other relevant examples covered here are the anisotropic quadratic costs and subspace elastic costs (already considered in Section 1.4 in the weakly log-concave regime).
2.6.2. If $\nu$ compactly supported
From the compactness of both marginals we are guaranteed that uniformly in $\operatorname{supp}(\rho)\times\operatorname{supp}(\nu)$ it holds
$h(c)\preceq\nabla_{2}^{2}c(x,y)\preceq H(c)\,,$
for some finite $h(c),\,H(c)$. Hence from (2.2)
$\nabla^{2}\psi^{n}_{\varepsilon}(y)\preceq-h(c)+\varepsilon^{-1}\max_{y\in\operatorname{supp}(\nu)}\left\|\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\nabla_{2}c(X,y))\right\|_{2}\preceq-h(c)+\varepsilon^{-1}\left\|\nabla_{2}c\right\|_{\mathrm{L}^{\infty}(\rho\times\nu)}^{2}\,,$
which yields
$\nabla^{2}(c(x,y)+\psi^{n}_{\varepsilon}(y))\preceq H(c)-h(c)+\varepsilon^{-1}\left\|\nabla_{2}c\right\|_{\mathrm{L}^{\infty}(\rho\times\nu)}^{2}\,.$
This, combined with Theorem 1.2, implies the exponential convergence of Sinkhorn’s algorithm and for
(2.31) $\varepsilon\leq\frac{\tau(H(c)-h(c))+\sqrt{\tau^{2}(H(c)-h(c))^{2}+4\tau\left\|\nabla_{2}c\right\|_{\mathrm{L}^{\infty}(\rho\times\nu)}^{2}}}{2}$
the rate is
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+\tau\varepsilon(H(c)-h(c))+\tau\left\|\nabla_{2}c\right\|_{\mathrm{L}^{\infty}(\rho\times\nu)}^{2}}\,.$
$\square$
2.7. Proof of Proposition 1.4
Let us briefly discuss how the gradient and the Hessian of any function $f\colon{\mathbb{S}^{d}}\to\mathbb{R}$ can be computed in the sphere Riemannian metric. For notations’ clarity, in the Riemannian setting we denote gradient and Hessian with $\nabla_{\mathbb{S}^{d}}$ and $\operatorname{Hess}_{\mathbb{S}^{d}}$, respectively. In view of this, let us fix $y\in{\mathbb{S}^{d}}$ as well as a tangent vector $v\in T_{y}\mathbb{S}^{d}\hookrightarrow\mathbb{R}^{d+1}$. This allows us to consider the constant-speed geodesic started at $y\in\mathbb{S}^{d}$ with velocity $v\in T_{y}\mathbb{S}^{d}$ as the curve $(\gamma_{t})_{t\in[0,1]}$ given by $\gamma_{t}=\exp_{y}(tv)$, explicitly given by
(2.32) $\gamma_{t}=\exp_{y}(tv)=\cos(t\|v\|)\,y+\sin(t\|v\|)\frac{v}{\|v\|}\,.$
To determine $\nabla_{\mathbb{S}^{d}}f(y)\in T_{y}{\mathbb{S}^{d}}$ and $\operatorname{Hess}_{{\mathbb{S}^{d}}}f(y):(T_{y}{\mathbb{S}^{d}})^{\otimes 2}\to\mathbb{R}$ it is enough to compute
(2.33) $(f\circ\gamma_{t})^{\prime}(0)=\langle\nabla_{\mathbb{S}^{d}}f(y),v\rangle_{\mathfrak{g}}\quad\text{ and }\quad(f\circ\gamma_{t})^{\prime\prime}(0)=\operatorname{Hess}_{{\mathbb{S}^{d}}}f(y)(v,v)\,.$
When considering the function $f(y)=c(x,y)=1-\langle x,y\rangle$, with $x\in{\mathbb{S}^{d}}$ fixed, it is not difficult to see that
$\nabla_{\mathbb{S}^{d}}c(x,\cdot)(y)=-(\operatorname{Id}-yy^{\intercal})x\qquad\textrm{and}\qquad\operatorname{Hess}_{\mathbb{S}^{d}}c(x,\cdot)(y)=\langle x,y\rangle\operatorname{Id}\,.$
This immediately entails the uniform bound
$-\mathfrak{g}\,\preceq\operatorname{Hess}_{\mathbb{S}^{d}}c(x,\cdot)\preceq\mathfrak{g}\,,$
with $\mathfrak{g}$ being the canonical Riemmanian metric of ${\mathbb{S}^{d}}$ given by the inclusion ${\mathbb{S}^{d}}\hookrightarrow\mathbb{R}^{d+1}$, from which we may deduce that on $T_{y}{\mathbb{S}^{d}}$ it holds
$\operatorname{Hess}_{\mathbb{S}^{d}}\psi^{n}_{\varepsilon}(y)\preceq\operatorname{Id}+\varepsilon^{-1}\,\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}\biggl{(}(\operatorname{Id}-yy^{\intercal})X\biggr{)}\preceq 1+\varepsilon^{-1}\,.$
Therefore the map $y\mapsto c(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave uniformly in $x\in{\mathbb{S}^{d}}$ and $n\in\mathbb{N}$ with $\Lambda=2+\varepsilon^{-1}$, which allows us to apply Theorem 1.2 and deduce the exponential convergence of Sinkhorn’s algorithm. Moreover, for any $\varepsilon\leq 1+\sqrt{1+\tau^{2}}$ the exponential convergence rate is equal to $1-\nicefrac{{\varepsilon^{2}}}{{\varepsilon^{2}+2\tau\varepsilon+\tau}}$. This proves the first part of Section 1.4.
Along the same lines, if we now consider $f(y)\coloneqq c_{\delta}(x,y)=\arccos(\delta\langle x,y\rangle)^{2}$, with $x\in{\mathbb{S}^{d}}$ fixed, and compute again (2.33), straightforward computations lead to
$\nabla_{\mathbb{S}^{d}}c_{\delta}(x,\cdot)(y)=-\frac{2\delta\,\arccos(\delta\langle x,y\rangle)}{\sqrt{1-\delta^{2}\langle x,y\rangle^{2}}}(\operatorname{Id}-yy^{\intercal})x$
and
(2.34) $\begin{split}\operatorname{Hess}_{\mathbb{S}^{d}}c_{\delta}(x,\cdot)(y)&=\frac{2\delta\,\arccos(\delta\langle x,y\rangle)}{\sqrt{1-\delta^{2}\langle x,y\rangle^{2}}}\langle x,y\rangle\operatorname{Id}\\ &\quad+2\delta^{2}\frac{1-\langle x,y\rangle^{2}}{1-\delta^{2}\langle x,y\rangle^{2}}\,\biggl{(}1-\delta\frac{\langle x,y\rangle\,\arccos(\delta\langle x,y\rangle)}{\sqrt{1-\delta^{2}\langle x,y\rangle^{2}}}\biggr{)}\,u\otimes u\,,\end{split}$
where $u$ is a unit-norm vector defined as $u\coloneqq\nicefrac{{(\operatorname{Id}-yy^{\intercal})x}}{{\left\|(\operatorname{Id}-yy^{\intercal})x\right\|}}\in T_{y}{\mathbb{S}^{d}}$. Let us also remark here that in the limit $\delta\uparrow 1$ the above quantities give the gradient and the Hessian of the squared distance function $y\mapsto{\sf d}^{2}(x,y)$ (in total concordance with the convergence $c_{\delta}\to{\sf d}^{2}$ as $\delta\uparrow 1$).
Given the above preliminary remarks, we are now ready to prove Section 1.4. Indeed, from (2.34) uniformly in $x\in{\mathbb{S}^{d}}$ we have
$-\frac{4\pi}{\sqrt{1-\delta^{2}}}\,\mathfrak{g}\,\preceq\operatorname{Hess}_{\mathbb{S}^{d}}c_{\delta}(x,\cdot)\preceq\biggl{(}2\delta^{2}+\frac{4\pi}{\sqrt{1-\delta^{2}}}\biggr{)}\,\mathfrak{g}\,.$
From these bounds it is also clear why we are restricting to costs $c_{\delta}$ with $\delta<1$.
As a consequence of the above discussion and (2.2) we have
$\operatorname{Hess}_{\mathbb{S}^{d}}\psi^{n}_{\varepsilon}(y)\preceq\frac{4\pi}{\sqrt{1-\delta^{2}}}+\varepsilon^{-1}\,\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}\biggl{(}\frac{2\delta\,\arccos(\delta\langle X,y\rangle)}{\sqrt{1-\delta^{2}\langle X,y\rangle^{2}}}(\operatorname{Id}-yy^{\intercal})X\biggr{)}$
$\preceq\frac{4\pi}{\sqrt{1-\delta^{2}}}+\varepsilon^{-1}\frac{8\pi^{2}}{1-\delta^{2}}\,.$
From this we immediately deduce that the map $y\mapsto c_{\delta}(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave uniformly in $x\in{\mathbb{S}^{d}}$ and $n\in\mathbb{N}$ with
$\Lambda=2\delta^{2}+\frac{8\pi}{\sqrt{1-\delta^{2}}}+\varepsilon^{-1}\frac{8\pi^{2}}{1-\delta^{2}}\,,$
which allows us to apply Theorem 1.2 and deduce the exponential convergence of Sinkhorn’s algorithm. Moreover, as soon as
(2.35) $\varepsilon\leq\tau\biggl{(}\delta^{2}+\frac{4\pi}{\sqrt{1-\delta^{2}}}\biggr{)}+\sqrt{\tau^{2}\biggl{(}\delta^{2}+\frac{4\pi}{\sqrt{1-\delta^{2}}}\biggr{)}^{2}+\frac{8\pi^{2}}{1-\delta^{2}}}$
the convergence rate is
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\tau\varepsilon\left(\delta^{2}+\frac{4\pi}{\sqrt{1-\delta^{2}}}\right)+\frac{8\pi^{2}\tau}{1-\delta^{2}}}\,.$
This concludes the proof of the second half of Section 1.4.
$\square$
2.8. Proof of Proposition 1.4
To establish this result, we first state an $\mathrm{L}^{\infty}$-bound for the gradient of the logarithm of the heat kernel.
[…]
Let $(M,\mathfrak{g})$ be a compact smooth Riemannian manifold. Then there is a constant $C$ depending on the Ricci curvature and on the dimension of $M$ only such that
(2.36) $t^{2}|\nabla_{M}\log\mathsf{p}_{t}(x,\cdot)(y)|^{2}\leq C(1+\kappa^{-}t)(1+t),\qquad\forall x,y\in M,\,\forall t>0\,,$
where $\kappa\in\mathbb{R}$ is a lower bound for the Ricci curvature.
If ${\rm Ric}_{M}\succeq 0$, then (2.36) has been proved in [39]. If instead ${\rm Ric}_{M}\succeq\kappa\mathfrak{g}$ with $\kappa<0$, then a slight adaptation of [37, Theorem 4.4] yields the same conclusion. For completeness, we postpone the proof to Appendix B. From this we immediately deduce that for any unit vector $v\in T_{y}M$
$\varepsilon^{-1}\,\mathrm{Cov}_{X\sim\pi^{n,n}(\cdot|y)}(\langle\nabla_{M}c_{\varepsilon}(X,\cdot)(y),v\rangle_{\mathfrak{g}})\leq C(1+\kappa^{-}(1+\varepsilon)+\varepsilon^{-1})\,.$
Next, we leverage a two-sided bound for the Hessian of the logarithm of the heat kernel obtained in [58], valid for $\varepsilon\leq 1$ (whence the necessity of this assumption). This result ensures the existence of a positive constant $C^{\prime}$ depending only on the Ricci curvature, the sectional curvatures, and on the dimension of $M$ such that
(2.37) $-C^{\prime}\left(1+\frac{{\rm diam}(M)^{2}}{\varepsilon}\right)\mathfrak{g}\preceq\operatorname{Hess}_{M}c_{\varepsilon}(x,\cdot)\preceq C^{\prime}\left(1+\frac{{\rm diam}(M)^{2}}{\varepsilon}\right)\mathfrak{g}\,.$
From this and (2.2) we deduce that
$\operatorname{Hess}_{M}\psi^{n}_{\varepsilon}\preceq\left(C^{\prime}(1+\varepsilon^{-1}{\rm diam}(M)^{2})+C(1+\kappa^{-}(1+\varepsilon)+\varepsilon^{-1})\right)\,\mathfrak{g}\,,$
so that the map $y\mapsto c_{\varepsilon}(x,y)+\psi^{n}_{\varepsilon}(y)$ is $\Lambda$-semiconcave uniformly in $x\in{\mathbb{S}^{d}}$ and $n\in\mathbb{N}$ with
$\Lambda=2C^{\prime}(1+\varepsilon^{-1}{\rm diam}(M)^{2})+C(1+\kappa^{-}(1+\varepsilon)+\varepsilon^{-1})\,.$
This allows us to apply Theorem 1.2 and deduce the exponential convergence of Sinkhorn’s algorithm. Moreover, as soon as
(2.38) $\varepsilon<\begin{cases}+\infty\quad&\text{ if }\tau C\kappa^{-}\geq 1\,,\\ \frac{2\tau C^{\prime}+\tau C+\kappa^{-}+\sqrt{(2\tau C^{\prime}+\tau C+\kappa^{-})^{2}+4\tau(1-\tau C\kappa^{-})(2C^{\prime}{\rm diam}(M)^{2}+C)}}{2(1-\tau C\kappa^{-})}\quad&\text{ otherwise,}\end{cases}$
the exponential convergence rate equals
$1-\frac{\varepsilon^{2}}{\varepsilon^{2}+2\tau C^{\prime}(\varepsilon+{\rm diam}(M)^{2})+\tau C(\varepsilon+\kappa^{-}\varepsilon(1+\varepsilon)+1)}\,.$
Let us finally comment the case $M$ has non-negative Ricci curvature. A careful look at [58] shows that the restriction to $\varepsilon\leq 1$ is needed to ensure the validity of Eq.(2.8) therein, namely
(2.39) $\frac{C_{0}}{{\rm vol}(B_{\sqrt{\varepsilon}}(x))}e^{-C_{1}\frac{{\rm diam}(M)^{2}}{\varepsilon}}\leq\mathsf{p}_{\varepsilon}(x,y)\leq\frac{C_{2}}{{\rm vol}(B_{\sqrt{\varepsilon}}(x))}\,,\qquad\forall x,y\in M,\,\forall\varepsilon\in(0,1]$
for appropriate constants $C_{0},C_{1},C_{2}$ depending only on the curvature $\kappa$ and the dimension $\dim(M)$, from which he deduces our (2.37). This is an adaptation to compact manifolds of Gaussian heat kernel estimates [45]. Indeed, on a manifold $M$ with ${\rm Ric}_{M}\succeq\kappa\mathfrak{g}$ it holds
$\frac{C_{0}}{{\rm vol}(B_{\sqrt{\varepsilon}}(x))}e^{-\frac{{\sf d}^{2}(x,y)}{3\varepsilon}-C_{\kappa}\varepsilon}\leq\mathsf{p}_{\varepsilon}(x,y)\leq\frac{C_{2}}{{\rm vol}(B_{\sqrt{\varepsilon}}(x))}e^{-\frac{{\sf d}^{2}(x,y)}{3\varepsilon}+C_{\kappa}\varepsilon}\,,\qquad\forall x,y\in M,\,\forall\varepsilon>0$
with $C_{\kappa}\geq 0$ solely depending on $\kappa$, and this implies both [58, Eq.(2.8)] and (2.39) if $\varepsilon\leq 1$. However, if ${\rm Ric}_{M}\succeq 0$, the constant $C_{\kappa}$ in the bound above can be chosen equal to 0, whence the validity of [58, Eq.(2.8)] and (2.39) for all $\varepsilon>0$. As a consequence, the upper and lower estimates on $\operatorname{Hess}_{M}c_{\varepsilon}(x,\cdot)$ stated at (2.37) holds true for all $\varepsilon>0$, whence the $\Lambda$-semiconcavity of $y\mapsto c_{\varepsilon}(x,y)+\psi^{n}_{\varepsilon}(y)$ with the same $\Lambda$ as above.
$\square$
[…]
To determine whether the condition $\tau C\kappa^{-}\geq 1$ in (2.38) is satisfied or not, we need a tractable expression for
$C=2\log\bigg{(}\frac{C_{2}C_{3}}{C_{0}}\bigg{)}\vee 4C_{1}{\rm diam}(M)^{2}$
namely for the constants $C_{0},C_{1},C_{2}$ appearing in the Gaussian heat kernel estimate (2.39) and on the doubling constant $C_{3}$, see (B.2). If ${\rm Ric}_{M}\succeq 0$ and we set $d=\dim(M)$, then $C_{3}=2^{d/2}$ while $C_{0},C_{1},C_{2}$ can be determined looking at the proof of [45, Thm. 3.1 and Cor. 3.1]. Indeed, (choosing for instance $\varepsilon=\delta=1$ in the aforementioned results) admissible values are $C_{1}=\frac{1}{3}$, $C_{0}=C_{2}^{-1}$ and
$C_{2}=\sup_{x>0}8^{d}(4x+1)^{d}e^{-\frac{x^{2}}{30}}\,.$
Appendix A
We write here a corollary of [19, Lemma 3.1] using the notation of this article.
[…]
Let $\varphi$ be such that for some $\alpha>0,L\geq 0$
$\kappa_{(\varepsilon^{-1}\varphi+U_{\rho})}(r)\geq\alpha\varepsilon^{-1}-r^{-1}f_{L}(r)\,\quad\forall r>0$
holds. Then we have
$\kappa_{\Phi^{\rho}_{0}(\varphi)}(r)\geq\frac{\varepsilon\alpha-1}{\varepsilon\alpha}-\frac{L}{\varepsilon\alpha^{2}}\quad\forall r>0.$
To connect the notation used here with that of [19] we observe that for all $\varphi$, we have that, upon setting $g=\varepsilon^{-1}\varphi+U_{\rho}$, we have $\Phi^{\rho}_{0}(\varphi)=U^{\varepsilon,g}_{0}$, where $[0,\varepsilon]\times\mathbb{R}^{d}\ni(s,x)\mapsto U^{\varepsilon,g}_{s}(x)$ is the solution of the Hamilton-Jacobi-Bellman equation
$\begin{cases}\partial_{s}\theta_{s}+\frac{1}{2}\Delta\theta_{s}-\frac{1}{2}|\nabla\theta_{s}|^{2}=0,\\ \theta_{\varepsilon}=g.\end{cases}$
Appendix B Proof of Lemma 2.8
As a starting point, recall Hamilton’s gradient estimate [38]: this ensures that if $u$ is a non-negative solution to the heat equation $\partial_{t}u=\Delta u$ with $u(0,\cdot)\in\mathrm{L}^{\infty}(M)$ and ${\rm Ric}_{M}\succeq\kappa\mathfrak{g}$ globally on $M$ for some $\kappa\in\mathbb{R}$, then
(B.1) $t|\nabla_{M}\log u(t,\cdot)|^{2}\leq(1+2\kappa^{-}t)\log\left(\frac{\|u(0,\cdot)\|_{\mathrm{L}^{\infty}}}{u(t,\cdot)}\right),\qquad\forall t>0\,.$
The desired bound (2.36) will follow by applying the inequality above to $u(s,y)=\mathsf{p}_{t+s}(x,y)$, where $t>0$ and $x\in M$ are fixed. To this end, we need to verify that $u(0,\cdot)=\mathsf{p}_{t}(x,\cdot)\in\mathrm{L}^{\infty}(M)$, since by the compactness of $M$ we already know that ${\rm Ric}\geq\kappa\operatorname{Id}$ for some $\kappa\in\mathbb{R}$. In particular, the fact that the Ricci curvature tensor is lower bounded grants (2.39), whence
$\begin{split}\|\mathsf{p}_{t}(x,\cdot)\|_{\mathrm{L}^{\infty}}&=\sup_{y\in M}\mathsf{p}_{t}(x,y)\leq\frac{C_{2}}{{\rm vol}(B_{\sqrt{t}}(x))}\,,\\ \inf_{x}\mathsf{p}_{2t}(x,y)&\geq\frac{C_{0}e^{-C_{1}\frac{{\rm diam}(M)^{2}}{t}}}{{\rm vol}(B_{\sqrt{2t}}(x))}>0\,.\end{split}$
By the Bishop–Gromov inequality (see for instance [50]), we know that for some constant $C_{3}>0$ it holds
(B.2) ${\rm vol}(B_{\sqrt{2t}}(x))\leq C_{3}{\rm vol}(B_{\sqrt{t}}(x))\,,\qquad\forall x\in M,\,\forall t>0\,,$
so that the above yields
$\frac{\|\mathsf{p}_{t}(x,\cdot)\|_{\mathrm{L}^{\infty}}}{\mathsf{p}_{2t}(x,y)}\leq\frac{C_{2}C_{3}}{C_{0}}e^{C_{1}\frac{{\rm diam}(M)^{2}}{t}}\,,\qquad\forall x,\,y\in M,\,\forall t\in(0,1]\,.$
We can now apply Hamilton’s gradient estimate with $u(s,y)=\mathsf{p}_{t+s}(x,y)$, as anticipated, and then set $s=t$, to get
$t|\nabla_{M}\log\mathsf{p}_{2t}(x,\cdot)(y)|^{2}$ $\leq(1+2\kappa^{-}t)\log\left(\frac{\|\mathsf{p}_{t}(x,\cdot)\|_{\mathrm{L}^{\infty}}}{\mathsf{p}_{2t}(x,y)}\right)$
$\leq(1+2\kappa^{-}t)\left(\log\bigg{(}\frac{C_{2}C_{3}}{C_{0}}\bigg{)}+C_{1}\frac{{\rm diam}(M)^{2}}{t}\right)\,,$
which is readily verified to be equivalent to (2.36) with $C=2\log\big{(}\frac{C_{2}C_{3}}{C_{0}}\big{)}\vee 4C_{1}{\rm diam}(M)^{2}$.
$\square$